### v1

# Data Analysis Prediction over Multiple Unseen Datasets: A Vector Embedding Approach

Abstract
The massive increase in the data volume and dataset availability for analysts compels researchers to focus on data content and select high-quality datasets to enhance the performance of analytics operators. While selecting the highest quality data for analysis highly increases task accuracy and efficiency, it is still a hard task, especially when the number of available inputs is very large. To address this issue, we propose a novel methodology that infers the outcome of analytics operators by creating a model from datasets similar to the queried one. Dataset similarity is performed via projecting each dataset to a vector embedding representation. The vectorization process is performed using our proposed deep learning model NumTabData2Vec, which takes a whole dataset and projects it into a lower vector embedding representation space. Through experimental evaluation, we compare the prediction performance and the execution time of our framework to another state-of-the-art modelling operator framework, illustrating that our approach predicts analytics outcomes accurately. Furthermore, our vectorization model can project different real-world scenarios to a lower vector embedding representation and distinguish between them.
Index Terms:
Analytics Modelling, Vector embeddings, Vector Similarity, Data Quality
I Introduction
Big data technologies daily face the rapid evolution in volume as well as variety and velocity of processed data [1]. Such big data characteristics routinely force analytics pipelines to underperform, requiring continuous maintenance and optimization. One major reason for this is bad quality of data [^1] https://tinyurl.com/de62sf48. Poor data quality leads to low data utilisation efficiency and even brings forth serious decision-making errors[2].
Data quality can be improved when focusing on the actual content of the data. Data-centric Artificial Intelligence (AI) [3] emphasises on the quality, context, and structure of the data to improve its quality, as well as the analytical or machine learning (ML) algorithmic performance. Understating the data context properties such as data features, origins, relevance, and potential biases plays a critical role in modelling more accurate and reliable models. Data-centric AI prioritises the process of refining and enriching datasets to make them more suitable for real-world applications. In a similar way, a lot of scientists make the case that improving data quality by focusing on the content is a crucial factor in achieving better results [3].
A plethora of available data sources and datasets in the organisations’ data centres and lakes pose a significant challenge: Deciding which datasets should be selected and passed into analytic workflows in order to produce more accurate results/predictions. While modern analytics workflows have evolved to incorporate diverse operators, the process of selecting the most suitable datasets using data-centric AI techniques to optimise a possibly complex computation is still under research [4]. When dataset selection is left to human experts, it can result in low prediction performance and can be time-consuming. Equally costly can be the process of exhaustively executing over all (or many) available dataset inputs, in order to identify high-quality inputs according to business standards.
In previous work [12], predicting the output of an analytics operator assuming a plethora of available input datasets was tackled via the creation of an all-pair similarity matrix, which, relative to the similarity function used, reflected the distance between datasets over a single characteristic (e.g., data distribution).
To improve the big data analysis and utilize modern AI methods and systems, data or vector embeddings have been suggested. Data embedding vectorization [14, 15] aims at projecting data from a high-dimensional representation space into a more compact, lower dimensional space. By extracting meaningful patterns from the data features using deep learning, data is projected to a lower representation space. Using this space, one can find the most similar data points among a dataset.
For the purpose of improving the accuracy of a modelled analytic operator (i.e., predict the outcome of a ML algorithm without actually executing it due to its cost), we propose a modelling framework that employs vector embeddings for dataset selection from a data lake store that contains a large number of available datasets. In order to predict an operator’s output for an “unseen” query dataset, our method considers only datasets that are qualitatively similar to the query, utilising a similarity search over the vector embeddings. The selection of similar datasets reduces the prediction error, as well as the cost to model the operator, under the assumption that realistic analytical operators perform similarly under similar inputs. Similarity search is thus executed on the dataset vector embedding representations in a low-dimensional space. Each vector embedding is generated using our proposed deep learning method, NumTabData2Vec, which takes an entire dataset as input and transforms it into a vector embedding representation space. While previous methods split the dataset into chunks and produced a vector for each chunk or used the dataset metadata information, we use the entire dataset utilising all data record values. Our framework thus models an operator using similar datasets to the queried one. An interesting feature of our methodology is the ability of the NumTabData2Vec to efficiently distinguish with detail between different datasets and record properties in the lower vector representation space. Furthermore, the same vector representation is used for the modelling of multiple analytics operators, offering enhanced flexibility and time-efficiency. Compared to similar previous work [12, 13], our work uses state-of-the-art data representation (vector embeddings) which are able to capture multiple data properties that can be used in order to assess similarity, namely record order, dataset size, data distribution, etc.
The main contributions of our work can be summarised as follows:
  • We introduce a framework for operator modelling in order to predict its outcome on an unseen tabular input dataset from a plethora of available ones. Our method uses dataset vector embedding representations to improve the prediction performance via selecting the most relevant datasets to base its prediction upon.
  • We develop a deep learning model architecture that transforms an entire tabular dataset of numerical values to a vector embedding representation.
  • We provide an experimental evaluation of our proposed methodology using multiple real-world scenarios and compare it directly to system Apollo [12, 13].
Our evaluation illustrates that our proposed methodology can model operators with low prediction error by adaptively selecting similar quality datasets for each model. Moreover, it is able to do so with considerable amortized speed-ups. The evaluation of NumTabData2Vec indicates that it accurately projects any available dataset into a vector embedding representation and accurately models different dataset properties within the representation space.
II Related Work
Prior efforts largely concentrate on boosting/scaling an algorithm’s performance utilizing only the size of data input (record number) without examining its quality. As such, we also discuss works that try to identify the most suitable data features for the optimisation of analytic task operators. Vectorising data to lower embedding representation is a modern method that helps in identifying significant features across data types and datasets. As vector embeddings extract important features from data, we discuss studies that used the feature representation of data tuples to improve ML model prediction.
II-A Data Quality
Big data applications aim to improve data quality by addressing various challenges. Dagger [8] enhances data quality by detecting errors in data pipelines using an SQL-like language and allowing code-handling for variables and arrays. ReClean [9] automates tabular data cleaning via reinforcement learning, optimizing tool selection for error repair. IterClean [10] employs a large language model (LLM) to iteratively clean data by labelling initial tuples and using error detection, verification, and repair. In [11], data tuple quality is measured using Shapley values from game theory, with Truncated Monte Carlo Shapley and Gradient Shapley methods estimating a tuple’s value to a learning algorithm. Apollo [12, 13] is a content-based method that predicts the outcome of an analytic operator by leveraging dataset similarity. It involves three steps: creating a similarity matrix, projecting datasets to a lower-dimensional space, and modelling the operator using a small random subset of datasets. Unlike Apollo, our approach selects the most relevant, high-quality datasets to model analytic operators, aiming to improve prediction performance. Moreover, modelling and similarity occur over the vector embeddings which are able to incorporate all dataset properties. In contrast, similarity functions used in [12] can only target a single property.
II-B Dataset Selection Inference
SOALA [26] selects optimal data features through online pairwise comparisons to maintain ML models over time. Its extension, Group-SOALA, introduces group maintenance to identify high-quality feature sets. In [27], the tf.data API framework enables the creation of ML pipelines focused on selecting relevant datasets and features to improve data quality. Similarly, our framework uses dataset vector embeddings to select the most suitable datasets for modelling analytic operators or ML models, enhancing prediction accuracy.
II-C Data Vectorization and Embeddings
The goal of data vectorization is to project high-dimensional data into a lower-dimensional vector space. Word2Vec [14] generates word embeddings using two neural network architectures, Continuous Bag of Words and Skip-Gram, leveraging word context. Graph2Vec [15] creates graph embeddings by dividing graphs into sub-graphs with a skip-graph model and aggregating their embeddings. ImageDataset2Vec [17] extracts meta-features from image datasets to generate embeddings, helping to select the most suitable classification algorithm. Dataset2Vec [18] uses meta-features and the DeepSet model to project datasets into embeddings and measure dataset similarity. Table2Vec [19] generates table embeddings by incorporating data features, metadata, and structural elements like captions and column headings. Inspired by these methods, we designed a model that generates vector representations of tabular datasets over their record data values, not their metadata.
Vector embeddings, which capture valuable information from data tuples, are widely used in classification tasks. TransTab [20] transforms data tuples into vector embeddings, applying transformer layers to encode features and predict classes, enhanced by supervised and self-supervised pretraining. FT-Transformer [21] and Res-Net architectures similarly use embeddings of categorical and numerical features, processed through transformer layers for class prediction. Tab-Transformer [22] embeds categorical features into vectors, normalizes continuous features, and combines them in a Multi-Layer Perceptron (MLP) to predict classes. Unlike these methods, which focus on tuple-level embeddings, our framework uses dataset-level vector embeddings to identify the most relevant datasets, improving analytic operator performance.
III Methodology
Figure 1: Pipeline framework architecture
In this section, we describe our proposed framework for modelling analytic operators over a large number of available input datasets. We also describe our approach for vectorizing tabular datasets, NumTabData2Vec.
III-A Framework Architecture
Consider a data lake store that contains a (possibly large) number $n$ of datasets $D=\left(d_{1},\ d_{2},\ d_{3},\ \ldots,\ d_{n}\right)$. Also, let us consider an analytics operator (e.g., a ML algorithm) $\Phi$ and an “unseen” dataset $D_{o}$. We assume that each $D_{i},1\leq i\leq n$ as well as $D_{o}$ consist of records with the same number of columns with numerical values only. Each dataset can, naturally, have a different number of records. Operator $\Phi$ consumes a single such dataset as input to produce a single numerical output: $\Phi:D_{i}\rightarrow\mathbb{R}$. Our goal is to predict $\Phi(D_{o}),\forall D_{o}$ with a minimal cost and prediction error. Our premise is that the output of an analytical operator on any unseen $D_{o}$ can be accurately modelled and runtime efficiency improved by selecting a small subset of $D$. This can be achieved by identifying the most similar datasets to $D$, denoted: $D_{o}$: $D_{r}\subseteq D_{o}$.
Datasets in $D_{r}$ closely match $D_{o}$ in their properties (e.g., order, distribution, and size to name a few). Previous work [12, 13] had to use separate similarity functions for each such property. In contrast, we leverage the embedding vectorization ($D_{z}$) to efficiently identify the most similar datasets using all dataset properties.
Figure 1 depicts the pipeline of our proposed framework. Datasets in $D$ are transformed into $k$-dimensional vectors via our NumTabData2Vec embedding scheme. Dataset vectorization needs no recomputation as the vectors are saved and can be reused for the rest of the framework steps. Each time a new dataset $D_{o}$ needs to be inferred relative to an analytics operator $\Phi$, its vector embedding is equally created. The datasets used for the creation of the model are selected via similarity search. Using different similarity functions, a small subset of datasets similar to $D_{o}$ are identified and used in order to create a model that predicts $\Phi(D_{o})$. With this approach, our framework is utilizing “right quality” data in its inference mechanism, with irrelevant and extraneous datasets being excluded from the modelling process.
The datasets are embedded by our NumTabData2Vec method. It takes an entire dataset as input and transforms it into a $k$-dimensional vector embedding representation space named $z$. The Vector embedding $z$, is a lower dimension representation of the dataset with the entire characteristics of the dataset being represented. Each dataset from $D$ is projected into such a $k$-dimensional vector embedding representation. The datasets are consequently represented by $D_{z}$ which contains all vector representations $\left(z_{1},\ z_{2},\ z_{3},\ \ldots,\ z_{n}\right)$. Dataset $D_{o}$ is also vectorized into $z_{o}$. Using the embedding representation $z_{o}$ and applying different similarity functions over the vector representations $D_{z}$, we may choose the most similar subset of $D$. The final step of the pipeline involves the operator modelling with any relevant method (e.g., Linear Regression, SVM, Multi-Layer Perceptron, etc.). This model can be used in order to infer the value of $\Phi(D_{o})$.
III-B NumTabData2Vec
Figure 2: NumTabData2Vec deep learning model architecture
This method transforms a dataset $D_{i}$ to a lower dimension representation $z$ by using only the numerical values of the dataset’s records and excluding metadata, column names, and filenames. We present a deep learning model architecture based on the variational autoencoder (VAE) [5] concept. For each dataset, data is projected from a high dimensional space to a lower vector embedding representation space $z$, with $1\times k$ dimension. The proposed method is thus defined as:
$NumTabData2Vec\left(D_{i}\right)\rightarrow z,$ (1)
where the model takes an $m\times n$ dimensional numerical dataset and projects it to a lower $k$-dimension space $z$ ($k>1$). We wish for our method to be generally applicable and able to use any scenario dataset that learns to project a vector embedding representation during the training phase. Also, we wish this method to be able to learn this vector embedding representation using a diverse amount of data during the training phase, and execute thereafter without the need of further training, fine-tuning, or online-training methods. Finally, we need our scheme to be able to quickly and precisely extract the vector for every dataset input as well as deal with varying dataset dimensions without requiring any modifications.
The deep learning model architecture is depicted in Figure 2, where a dataset $D_{i}$ with dimensions $m\times n$ passes through the encoder and is projected into a vector embedding representation $z$. The dataset has to be recreated since the decoder uses the same layers as the encoder but in the opposite direction. The learning of the vector representation $z$ is using the probabilistic encoder $q_{\phi}\left(z|x\right)$ distribution and the probabilistic decoder $p_{\theta}\left(x|z\right)$ distribution utilising the Kullback–Leibler (KL) divergence [6]. To achieve that, the following condition:
$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)$ (2)
of Kullback–Leibler (KL) divergence must be minimised. To learn the new vector representation $z$, the dataset must be reconstructed back from $z$ to their input format using the decoder, to verify if the vector representation is on a lower dimension in terms of the input dataset. This is the reconstruction loss. The loss function can be described as:
$\mathcal{L}_{\theta,\phi}\left(x\right)=\mathbb{E}_{q_{\phi}\left(z|x\right)}\left(\log{\left(p_{\theta}\left(x|z\right)\right)}\right)-D_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)$
This loss function is called evidences lower bound (ELBO). While the KL divergence is minimised to learn the vector embedding representation $z$, the ELBO is maximized so the condition,
$argmax\mathcal{L}_{\theta,\phi}\left(x\right)$ (4)
must be satisfied. The Decoder $p_{\theta}\left(x|z\right)$ is only used during the training phase to teach the encoder how to project the vector embedding representation $z$.
The decoder takes as an input the dataset $D_{i}$, first extracts the feature embeddings from the input layer and passes them into the Transformer layers. For the transformer layer we are using the pre-LN Transformer layer [7] instead of the traditional post-LN transformer layer where the normalisation layer is employed inside the residual connection and before the prediction of the feed-forward layer. Our model design takes into account and uses the Pre-LN Transformer layer as it demonstrates improved learning outcomes with shorter training times [7]. Following that, the transformed embedding space is projected onto a probabilistic distribution vector space $z$ using the mean ($\mu$) and standard deviation ($\sigma$). Then the transformed embedding space using $\mu$ and $\sigma$ are projected into a probabilistic distribution vector space $z$. The vector embedding space $z$ is a lower-dimensional representation of the dataset $D_{i}$ that preserves all the essential information needed to describe the dataset within the space $z$. The higher the $k$-dimension of the $z$ vector embedding representation, the higher and more accurate is this representation, as more features are extracted from the dataset [18].
III-C Dataset Selection
Algorithm 1 K-Means Clustering Algorithm for dataset selection
1:Data Lake Vectors $\mathbf{V}=\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{n}\}$, $D_{o}$ vector $v_{o}$, range of clusters $S=(2,\dots,p)$
2:Initialize the number of cluster $s$ using Silhouette score: $s=SilhouetteScore(V,S)$
3:Initialize the $s$ cluster centroids $\mathbf{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{s}\}$ randomly from the vectors $\mathbf{V}$.
4:repeat
5:     Assignment Step:
6:     for each vector $\mathbf{v}_{i}\in\mathbf{V}$ do
7:         Assign $\mathbf{v}_{i}$ to the nearest centroid based on Euclidean distance:
Assign $\text{Assign }\mathbf{v}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{v}_{i}-\mathbf{c}_{j}\|^{2}$
8:     end for
9:     Update Step:
10:     for each centroid $\mathbf{c}_{j}$ do
11:         Update $\mathbf{c}_{j}$ as the mean of all vectors assigned to cluster $j$:
$\mathbf{c}_{j}=\frac{1}{|\{\mathbf{v}_{i}\in C_{j}\}|}\sum_{\mathbf{v}_{i}\in C_{j}}\mathbf{v}_{i}$
12:     end for
13:until Centroids $\mathbf{C}$ do not change significantly
14:Save the cluster model K-Means
15:Find in which cluster the vector $v_{o}$ belongs,     $c=\textbf{K-Means}(v_{o})$
16:Find which datasets $D_{r}$ are belongs to cluster $c$
17:Return Datasets $D_{r}$
The selection of the most similar datasets has been implemented using three different approaches. Different similarity functions are easily plugged into our pipeline. The first method uses the cosine similarity to determine how similar two datasets are in their vector embedding space $z$. Cosine similarity calculates the angle between the two vectors and a measure of similarity that is independent of the vector magnitudes. The higher its value, the most similar are the two datasets. The following equation describes the cosine similarity between vectors $A$ and $B$:
$\cos\left(\theta\right)=\dfrac{A\cdot B}{\left\|A\right\|_{2}\left\|B\right\|_{2}}=\dfrac{\sum_{i=1}^{n}A_{i}B_{i}}{\sqrt{\sum_{i=1}^{n}A^{2}_{i}}\sqrt{\sum_{i=1}^{n}B^{2}_{i}}}$ (5)
The alternative method involves calculating the Euclidean distance between the two vectors in order to determine how similar the two datasets are. Euclidean distance is particularly useful for capturing the geometric closeness between vectors and finding the most relevant datasets based on the organisation’s requirements. The Euclidean distance between vectors $\mathbf{u},\mathbf{v}$ is defined as:
$d(\mathbf{u},\mathbf{v})=\sqrt{\sum_{i=1}^{n}(u_{i}-v_{i})^{2}}$ (6)
The smaller the distance value the most similar are the datasets. Dataset selection using cosine similarity or euclidean distance selects a percentage of $\lambda,\lambda>0$ of the closest datasets to dataset $D_{o}$. The third approach involves utilising the K-Means [23, 24] clustering technique to choose the most relevant datasets. The datasets from $D$ are divided into $s$ ($s>1$) separate clusters, where datasets with similar features are grouped together in the same cluster based on similarity equations. The silhouette scores [25] are used to determine the number of clusters $s$ for each possible value of $S$. This is done by the following equation:
$SilhouetteScore(v_{i})=\frac{b(v_{i})-a(v_{i})}{\max(a(v_{i}),b(v_{i}))}$ (7)
where for each vector $v_{i}$ finds the mean intra-cluster distance ($a(v_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(v_{i})$) is the minimum average distance with the other vectors in the different cluster. Silhouette Score ranges from $-1$ to $1$ and the higher value defines the best $s$ number for clusters.
Algorithm 1, shown how k-means clustering is created and how the most relevant datasets $D_{r}$ are selected based on the target dataset $D_{o}$. Using the most optimal $s$ value with the highest Silhouette score, K-means clustering is performed on the vector representations $V$ of each dataset, grouping datasets with same features into the same cluster. Next, the algorithm uses the vector $v_{o}$ of dataset $D_{o}$ to find the closest cluster centroid. All datasets in that cluster are defined as the relevant datasets $D_{r}$. These datasets are then used to model the analytics operator at hand.
IV Evaluation
We compared our framework with Apollo [12, 13], which models analytic operators using data content. Two loss functions to measure prediction accuracy are employed: root-mean-square error (RMSE) and mean absolute error (MAE). RMSE is sensitive to outliers, while MAE is not; conversely, RMSE accounts for error direction, which MAE cannot. Speedup metrics are also used to evaluate how efficiently our framework models operator $\Phi$. Specifically, Speedup and Amortized Speedup measure the time required to approximate each operator versus exhaustively executing them on all datasets. Speedup is defined as $\frac{T{{}^{(i)}_{op}}}{T{{}^{(i)}_{SimOp}+T_{vec}+T_{sim}+T_{pred}}}$, where $T{{}^{(i)}_{op}}$ is the time to execute operator $i$ on all datasets, $T{{}^{(i)}_{SimOp}}$ is the time to model the operator with datasets from similarity search, $T_{vec}$ is the vector embedding computation time, $T_{sim}$ is the similarity search time, and $T_{pred}$ is the prediction time for $D_{o}$. Amortized speedup includes dataset vectorization, performed once per data lake for multiple operators (in our case two operators). We also evaluate our dataset vectorization model, NumTabData2Vec, which projects datasets into vector embedding space $z$. Three versions were built with vector sizes of $100$, $200$, and $300$, each featuring eight transformer layers. The models were trained for 50 epochs on four NVIDIA A10 GPUs in parallel.
IV-A Evaluation Setup
Our framework is deployed over an AWS EC2 virtual machine server running with 48 VCPUs of AMD EPYC 7R32 processors at 2.40GHz, and four A10s GPUs with 24GB of memory each, $192GB$ of RAM memory, and $2TB$ of storage, running over Ubuntu 24.4 LTS. Our code is written in Python (v.3.9.1) and PyTorch modules (v.2.4.0). Apollo was deployed in a virtual machine with 8 VCPUs Intel Xeon E5-2630 @ 2.30GHz, $64GB$ of RAM memory, and $250GB$ of storage, running Ubuntu 24.4 LTS like in their experimental evaluation.
IV-B Datasets
TABLE I: Dataset properties for experimental evaluation
Dataset Name # Files # Tuples # Columns
Household Power
Consumption [28] $401$ $2051$ 7
Adult [29] $100$ $228$ 14
Stocks [30] $508$ $1959-13$ 7
Weather [31] $49$ $516$ 7
We evaluated our framework using four diverse datasets to represent real-world scenarios. Table I summarizes these datasets and their attributes. The vectorization module, NumTabData2Vec, was trained on data separate from the experimental evaluation data, split $60\%$ for training and $40\%$ for testing. The Household Power Consumption (HPC) dataset [28] contains electric power usage measurements from a household in Sceaux, France. It includes $401$ datasets, each with $2051$ tuples and seven features recorded at one-minute intervals. The Adult dataset [29], commonly used for binary classification, predicts whether an individual earns more or less than $50K$ annually. It comprises $100$ datasets, each with $228$ individuals and various socio-economic features. The Stock Market dataset [30] includes daily NASDAQ stock prices obtained from Yahoo Finance, with $508$ datasets. Each dataset contains $13$ to $1959$ tuples, each describing seven feature attributes. The Weather dataset [31] provides hourly weather measurements from $36$ U.S. cities between $2012$ and $2017$, split into $49$ datasets, each with $516$ tuples and seven features.
(a) Linear Regression RMSE error loss
(b) Linear Regression MAE error loss
(c) MLP for Regression RMSE error loss
(d) MLP for Regression MAE error loss
Figure 3: Household power consumption dataset prediction error loss
Our framework was evaluated by registering the accuracy of predicting the output of various ML operators over multiple datasets in $D$ without actually executing the operator on them. To evaluate our scheme and its parameters, we use all four datasets, ranging the size of the produced vectors as well as the similarity functions used. We project all datasets into $k$-dimensional spaces with varying vector dimensions ($100$, $200$, and $300$). For each dataset in Table I, we model different operators: For the regression datasets (Household Power Consumption and Stock Market), we model Linear Regression (LR) and Multi-Layer Perceptron (MLP) operators; for the classification datasets (Weather and Adult), we model the Support Vector Machine (SVM) and MLP classifier operators. Each experiment has been executed $10$ times and we report the average of the error loss, as well as the speedup.
(a) Linear Regression RMSE error loss
(b) Linear Regression MAE error loss
(c) MLP for Regression RMSE error loss
(d) MLP for Regression MAE error loss
Figure 4: Stock market dataset prediction error loss
IV-C Evaluation Results
Figures 3, 4, 5, and 6 present the evaluation results for each method, comparing the performance of different similarity search techniques across various vector embedding representation spaces. The red (with hatches), brown, and blue bars correspond to vector embeddings of size 100, 200, and 300 respectively. In each sub-figure, the y-axis represents the error loss value, while the x-axis displays the similarity search method applied over the vector embeddings. Figures 3 and 4 show the results for the Stock market and Household power consumption datasets, where the bottom sub-figure demonstrates the MLP regression model, and the top sub-figure presents the LR model. Figures 5 and 6 depict the evaluation results for the Weather and Adult datasets. In these Figures, the top sub-figure shows the SVM with SGD results, while the bottom sub-figure shows the MLP classifier. The left sub-figures in all Figures use the RMSE loss function, whereas the right sub-figures use the MAE loss function.
Figure 3, we show, for the HPC dataset, shows as increase the vector dimension size there is slightly lower prediction error for all the operator modelling. While for different similarity methods did not result in any significant differences in the prediction error loss for all the operator modelling. This suggests that, regardless the similarity selection method, our framework effectively selects the most optimal subset of data to improve model predictions on the unseen input dataset $D_{o}$. Additionally, we observe higher error loss with a vector size of 100, which can be attributed to the reduced representation capacity of lower-dimensional vectors. This limitation results in fewer “right” datasets being selected.
For the stock market dataset, Figure 4 depicts that a vector embedding representation of size $300$ models more accurate operators, with cosine similarity performing best in the similarity search and modelling the most optimal operator. However, due to the inherent volatility in Stock market data from different days, all models in the stock market dataset experiments exhibit high loss values.
In the weather dataset, the SVM operator results from sub-figures 5(a) and 5(b) show that using $300$ vectors in the representation space consistently led to more accurate operator models across all similarity methods. Specifically, cosine similarity in combination with the $300$-dimensional vector embedding reduced the error rate in operator predictions, demonstrating that projecting datasets into this representation space and applying cosine similarity improves the prediction accuracy on the modelled operator. For the MLP classifier from sub-figures 5(c) and 5(d), the results illustrate that using vector embeddings of size $200$ and K-Means clustering produced the most accurate MLP classifier operators.
(a) SVM with SGD RMSE error loss
(b) SVM with SGD MAE error loss
(c) MLP RMSE error loss
(d) MLP MAE error loss
Figure 5: Weather dataset prediction error loss
On the other hand, the Adult dataset shows the lowest error rates, with error loss values consistently below $0.5$ across all vector embedding dimensions and similarity search methods (see Figure 6). The Adult dataset, besides exhibiting a high number of rows, also has a higher number of columns, which demonstrates that our framework performs consistently well even with larger datasets. Additionally, we observe that the lowest prediction error across all datasets occurs when using higher-dimensional vector embeddings. With a trade-off between accuracy and execution time as the difference to generate all data lake available datasets vector embedding representation between $100$ and $300$ size dimension in the vector representation space to be less than $60$ seconds. This confirms that a higher number of vector dimensions leads to more accurate predictions, consistent with findings in previous research [14].
(a) SVM with SGD RMSE error loss
(b) SVM with SGD MAE error loss
(c) MLP RMSE error loss
(d) MLP MAE error loss
Figure 6: Adult dataset prediction error loss
We conducted an experimental evaluation using the Sampling Ratio (SR) approach, similar to Apollo [12], but employed neural networks built from the vector embeddings of each dataset. The SR approach involves a unified random selection of $l\%$ datasets from the vector representation space, using this subset to construct a neural network for predicting operator outputs. We tested SR values of $0.1$, $0.2$, and $0.4$, as well as vector embedding dimensions of $100$, $200$, and $300$, across all datasets. Figure 7 presents the sampling ratio results for the Adult dataset using MLP (sub-figure 7(a)) and for the Weather dataset using LR (sub-figure 7(b)). In each sub-figure the y-axis represents the RMSE prediction error loss while the x-axis denotes the vector dimension
(a) Adult Dataset MLP Operator RMSE error loss
(b) HPC dataset LR Operator RMSE error loss
Figure 7: Sampling Ratio prediction results
Both experiments demonstrate that as the vector embedding dimension increases, coupled with a larger sampling ratio (SR) value, there is a slight decrease in the prediction error loss. This improvement occurs because higher-dimensional vector embeddings provide a more accurate representation of the datasets in k-dimensions, with better dataset selection leading to enhanced prediction accuracy. Comparing the SR approach to our similarity search method for the HPC dataset, the SR approach was approximately $15\%$ less accurate in operator prediction across all vector embedding dimensions. A similar trend was observed in the Weather dataset. However, the Stock dataset exhibited a much larger discrepancy, with the SR approach performing about $70\%$ worse in prediction accuracy across all vector embedding dimensions. Likewise, in the Adult dataset, the SR approach delivered the poorest performance, with nearly $90\%$ lower prediction accuracy compared to the similarity search methods.
TABLE II: Evaluation results of our framework exported analytic operator with lowest prediction error in comparison with Apollo
Dataset Name Method Operator RMSE MAE Speedup Amortized Speedup
Household Power Consumption $300$V Cosine LR $\mathbf{6.61}$ $\mathbf{5.42}$ $0.0017$ $\mathbf{1.99}$
$300$V SR-$0.2$ LR $7.77$ $6.66$ $0.0018$ $1.42$
Apollo-SR $0.1$ LR $2968.01$ $2352.55$ $\mathbf{0.015}$ $0.024$
Apollo-SR $0.2$ LR $2811.49$ $2229.50$ $0.015$ $0.024$
$300$V K-Means MLP Regr. $\mathbf{6.70}$ $\mathbf{3.38}$ $0.9249$ $\mathbf{1.99}$
Apollo-SR $0.1$ MLP Regr. $3322.05$ $2606.99$ $2.38$ $1.74$
Apollo-SR $0.2$ MLP Regr. $3850.01$ $2609.36$ $\mathbf{2.38}$ $1.74$
Stock $300$V Cosine LR $306382.28$ $125335.65$ $0.00085$ $\mathbf{1.91}$
$300$V SR-$0.4$ LR $21861625.91$ $5674215.265$ $0.00087$ $0.33$
Apollo-SR $0.1$ LR $\mathbf{153665.92}$ $\mathbf{118236.48}$ $\mathbf{0.00093}$ $0.00096$
Apollo-SR $0.2$ LR $166844.95$ $133306.68$ $0.00093$ $0.00096$
$300$V Cosine MLP Regr. $\mathbf{140236.47}$ $\mathbf{123571.12}$ $0.63$ $\mathbf{1.91}$
Apollo-SR $0.1$ MLP Regr. $175150.82$ $145123.09$ $\mathbf{0.93}$ $0.96$
Apollo-SR $0.2$ MLP Regr. $174390.81$ $146338.73$ $0.93$ $0.96$
Weather $300$V Cosine SVM SGD $\mathbf{14.13}$ $\mathbf{7.63}$ $1.06$ $\mathbf{22.8}$
Apollo-SR $0.1$ SVM $69.51$ $25.52$ $\mathbf{2.10}$ $1.16$
Apollo-SR $0.2$ SVM $68.70$ $22.81$ $2.10$ $1.16$
$200$V Cosine MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
$200$V SR-$0.4$ MLP $15.95$ $13.31$ $1.02$ $1.77$
Apollo-SR $0.1$ MLP $69.62$ $23.10$ $\mathbf{1.34}$ $1.14$
Apollo-SR $0.2$ MLP $673.56$ $\mathbf{84.70}$ $1.32$ $1.14$
Adult $300$V Cosine SVM SGD $\mathbf{0.36}$ $\mathbf{0.2}$ $0.37$ $\mathbf{2.78}$
Apollo-SR $0.1$ SVM $68.32$ $22.95$ $\mathbf{0.75}$ $0.85$
Apollo-SR $0.2$ SVM $68.88$ $22.88$ $0.74$ $0.85$
$300$V K-Means MLP $\mathbf{0.36}$ $\mathbf{0.19}$ $0.30$ $2.78$
$300$V SR-$0.2$ MLP $6.01$ $6.00$ $0.54$ $\mathbf{3.54}$
Apollo-SR $0.1$ MLP $71.11$ $26.51$ $\mathbf{1.07}$ $1.31$
Apollo-SR $0.2$ MLP $70.16$ $25.74$ $1.05$ $1.31$
Table II compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$. Methods $100$V, $200$V, and $300$V denote vector embedding dimensions. The lowest prediction errors align with our pipeline’s similarity search method. Apollo outperforms our framework on the Stock dataset for the LR analytic operator at SR equals with $0.1$ (with $50\%$ and $6\%$ improvements for RMSE and MAE, respectively). However, our framework excels with the MLP regression operator, improving RMSE and MAE by $20\%$ and $17\%$, respectively. The LR operator’s performance gap on the Stock dataset is minor. For other datasets, our framework consistently surpasses Apollo across different SR values. This demonstrates the effectiveness of our similarity search approach, which enhances data quality and reduces $\Phi$ prediction errors by identifying relevant datasets $D_{r}$ from the data lake directory $D$. The Adult dataset also highlights our framework’s advantage with increasing feature dimensions. Although Apollo achieves better raw speedup due to the higher complexity of our framework’s vectorization step, our framework outperforms it in amortized speedup. By excluding the reusable vectorization process, it achieves speed gains of $10\%$ to $60\%$ for most operators. The SR approach, leveraging vector embedding representations, enhances operator prediction compared to Apollo and achieves greater amortized speedup. However, the similarity search method outperforms both Apollo and the SR approach in prediction accuracy and amortized speedup, establishing its clear superiority across most datasets and operator scenarios.
IV-D NumTabData2Vec Evaluation Results
Figure 8: Vector representation for each dataset from NumTabData2Vec
TABLE III: Similarity between vectors of different datasets scenarios
Model Name Similarity
NumTabData2Vec $100$ Vector size $0.54$
NumTabData2Vec $200$ Vector size $0.18$
NumTabData2Vec $300$ Vector size $0.16$
Our proposed model, NumTabData2Vec, was evaluated to determine its ability to distinguish dataset scenarios based on qualitative differences. For each scenario, $n$ random datasets were selected, and their vector embeddings averaged to represent the scenario. These embeddings, initially in a 200-dimensional space, were projected into 3D using PCA and are shown in Figure 8. The figure illustrates that each dataset scenario occupies a distinct space, with minimal overlap or clustering. This demonstrates that NumTabData2Vec effectively distinguishes datasets, outperforming prior methods like Word2Vec and Graph2Vec [14, 15], which achieved similar accuracy but on different data types (e.g., words, graphs) rather than entire datasets. Table III further highlights the average cosine similarity between dataset embeddings, showing greater dissimilarity as vector dimensions increase. However, results suggest that dimensions between $100$ and $300$ are sufficient for accurate distinction, avoiding the need for larger vector sizes.
Figure 9: Synthetic data vector embedding representation
To evaluate NumTabData2Vec’s ability to distinguish datasets with varying row and column counts, we generated synthetic numerical tabular datasets of different dimensions and vectorized them. Figure 9 shows datasets with columns ranging from three to thirty and rows from ten to one thousand, projected from a $200$-dimensional space to 2D using PCA. Each bullet caption c and r corresponds to the columns and rows of the dataset, respectively. Datasets with the same number of columns cluster closely in the representation space, and a similar pattern is observed for datasets with the same number of rows. These results indicate that our method effectively distinguishes datasets based on size during vectorization.
TABLE IV: NumTabData2Vec execution time for different dataset dimensions and different vector sizes
# of columns # of rows $50$ Vectors Execution time $100$ Vectors Execution time $200$ Vectors Execution time $3$ $100$ $0.0004$ sec $0.00042$ sec $0.00051$ sec $3$ $500$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1000$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1500$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $3$ $1800$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $10$ $100$ $0.0004$ sec $0.0004$ sec $0.00057$ sec $10$ $500$ $0.00039$ sec $0.0004$ sec $0.00051$ sec $10$ $1000$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $10$ $1500$ $0.00041$ sec $0.00042$ sec $0.00055$ sec $10$ $1800$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $20$ $100$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $500$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $1000$ $0.00042$ sec $0.00043$ sec $0.00052$ sec $20$ $1500$ $0.00043$ sec $0.00044$ sec $0.00054$ sec $20$ $1800$ $0.00044$ sec $0.00044$ sec $0.00054$ sec
To evaluate how dataset dimensions affect the execution time of NumTabData2Vec, we created synthetic datasets with varying numbers of rows ($100$, $500$, $1000$, $1500$, and $1800$) and columns ($3$, $10$, and $20$). These datasets were vectorized into different dimensions, and the execution times were recorded. Table IV shows that increasing the k-dimension requires approximately $20\%$ more time to generate the vector embeddings. This is expected, as a higher k-dimension involves more hyperparameters, which naturally increases computation time.
Interestingly, varying the number of columns did not significantly impact execution time. However, increasing the number of rows resulted in approximately $5\%$ additional execution time. This is because larger datasets require the extraction of more features, which has a modest impact on the model’s execution time.
Figure 10: HPC Dataset vector embedding representation with addition of Noise
To evaluate NumTabData2Vec’s ability to distinguish datasets based on different properties like distribution and order, we introduced Gaussian noise to random $l\%$ of data tuples in an HPC dataset. Figure 10 visualises the original and noise-modified datasets, projected from a 200-dimensional space to 2D using PCA. Each bullet caption g denotes the percentage of Gaussian noise added in the dataset. As noise increases, the representation space shifts further from the original dataset, indicating that NumTabData2Vec effectively captures distribution differences. Additionally, since the HPC dataset has an inherent order, the model’s sensitivity to noise demonstrates its ability to distinguish datasets based on ordering as well.
Figure 11: HPC Dataset vector embedding representation with addition of Noise in the first column
To evaluate how fine-grained as distinction can be, we introduced noise into a single column and repeated the previous experiment, with the difference being that noise was added exclusively to the first column. Figure 11 visualizes the dataset’s 2D vector space. The amount of Gaussian noise added to the dataset’s first column is indicated by g in the bullet caption. The results show that as more noise is introduced to the column, the vector representation moves further away from the original dataset. In contrast to the previous experiment shown in Figure 10, the noisy dataset’s representation stays closest to the original when only a single column is modified. Also in this experiment the dataset points in the 2-dimension are more grouped between them instead the previous experiment.
V Limitations
There exist a number of limitations in our work as we described it. In this section we briefly highlight them. Firstly, our input datasets comprise records of specific size and type (numerical). This currently excludes data with textual and categorical attributes, or tables with varying number of features inside a set of datasets. Secondly, we currently consider single-input and single-output operator modelling. Finally, our proposed NumData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples. This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available).
VI Conclusion
In this paper, we presented a novel framework for the modelling of an analytic operator (such as a ML algorithm) when a large number of input data is available and thus no brute-force execution can be performed. We propose a deep learning model, NumData2Vec, which transforms a dataset to a lower $k$-dimensional representation space $z$. Our framework produces vector embeddings for the input datasets using NumData2Vec and performs a similarity search to identify the most relevant subset of datasets for any unseen input. By modelling the analytic operator based on this selected subset, we are able to accurately predict its output on any given input dataset. In practice, we demonstrated that our framework can accurately model various common algorithms and compared favourably against a similar recent framework [12], in both accuracy and speedup. Furthermore, we showed that NumData2Vec can create different vector representations for datasets from different scenarios. We also demonstrated that NumData2Vec can effectively detect when noise is introduced into a dataset.
VII ACKNOWLEDGMENTS
This work is partially supported by the project RELAX-DN, funded by the European Union under Horizon Europe 2021-2027 Framework Programme Grant Agreement number 101072456.

### v2

@@
 Abstract
-The massive increase in the data volume and dataset availability for analysts compels researchers to focus on data content and select high-quality datasets to enhance the performance of analytics operators.
-While selecting the highest quality data for analysis highly increases task accuracy and efficiency, it is still a hard task, especially when the number of available inputs is very large.
-To address this issue, we propose a novel methodology that infers the outcome of analytics operators by creating a model from datasets similar to the queried one.
-Dataset similarity is performed via projecting each dataset to a vector embedding representation.
-The vectorization process is performed using our proposed deep learning model NumTabData2Vec, which takes a whole dataset and projects it into a lower vector embedding representation space.
-Through experimental evaluation, we compare the prediction performance and the execution time of our framework to another state-of-the-art modelling operator framework, illustrating that our approach predicts analytics outcomes accurately.
-Furthermore, our vectorization model can project different real-world scenarios to a lower vector embedding representation and distinguish between them.
+The rapid growth in data volume and the increasing availability of datasets present analysts with the challenge of selecting high-quality data to optimize the performance of analytics operations.
+While choosing the best-quality data significantly improves accuracy and efficiency, it remains a complex task, particularly when dealing with a vast number of inputs.
+To tackle this challenge, we introduce a novel approach that predicts analytics outcomes by building a model based on datasets similar to the target dataset.
+Our method determines dataset similarity by mapping each dataset into a vector embedding representation.
+This transformation is achieved through our deep learning model, NumTabData2Vec, which processes entire datasets and projects them into a lower-dimensional vector space.
+Through experimental evaluation, we compare our framework’s prediction accuracy and execution time against a leading state-of-the-art modelling operator, demonstrating that our approach effectively predicts analytics outcomes.
+Additionally, our vectorization model can capture and differentiate diverse real-world scenarios by representing them in a compact vector space.
 Index Terms:
@@
 Algorithm 1 K-Means Clustering Algorithm for dataset selection
-1:Data Lake Vectors $\mathbf{V}=\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{n}\}$, $D_{o}$ vector $v_{o}$, range of clusters $S=(2,\dots,p)$
-2:Initialize the number of cluster $s$ using Silhouette score: $s=SilhouetteScore(V,S)$
-3:Initialize the $s$ cluster centroids $\mathbf{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{s}\}$ randomly from the vectors $\mathbf{V}$.
+1:Vectors $\mathbf{Z}=\{\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}$, $D_{o}$ vector $z_{o}$, range of clusters $S=(2,\dots,p)$, Maximum size of cluster $max_{s}$, Minimum size of cluster $min_{s}$
+2:Initialize the number of cluster $s$ using Silhouette score: $s=SilhouetteScore(Z,S)$
+3:Initialize the $s$ cluster centroids $\mathbf{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{s}\}$ randomly from the vectors $\mathbf{Z}$.
 4:repeat
 5:     Assignment Step:
-6:     for each vector $\mathbf{v}_{i}\in\mathbf{V}$ do
-7:         Assign $\mathbf{v}_{i}$ to the nearest centroid based on Euclidean distance:
-Assign $\text{Assign }\mathbf{v}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{v}_{i}-\mathbf{c}_{j}\|^{2}$
+6:     for each vector $\mathbf{z}_{i}\in\mathbf{Z}$ do
+7:         Assign $\mathbf{z}_{i}$ to the nearest centroid based on Euclidean distance:
+Assign $\text{Assign }\mathbf{z}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|^{2}$
 8:     end for
@@
 11:         Update $\mathbf{c}_{j}$ as the mean of all vectors assigned to cluster $j$:
-$\mathbf{c}_{j}=\frac{1}{|\{\mathbf{v}_{i}\in C_{j}\}|}\sum_{\mathbf{v}_{i}\in C_{j}}\mathbf{v}_{i}$
+$\mathbf{c}_{j}=\frac{1}{|\{\mathbf{z}_{i}\in C_{j}\}|}\sum_{\mathbf{z}_{i}\in C_{j}}\mathbf{v}_{i}$
 12:     end for
@@
 14:Save the cluster model K-Means
-15:Find in which cluster the vector $v_{o}$ belongs,     $c=\textbf{K-Means}(v_{o})$
+15:Find in which cluster the vector $z$ belongs, $c=\textbf{K-Means}(z)$
 16:Find which datasets $D_{r}$ are belongs to cluster $c$
-17:Return Datasets $D_{r}$
+17:Check the number of datasets in $D_{r}$ and update it if it does not meet the $min_{s}$ and $max_{s}$.
+18:Return Datasets $D_{r}$
 The selection of the most similar datasets has been implemented using three different approaches.
@@
 These datasets are then used to model the analytics operator at hand.
+However, we defined a maximum and minimum size for $D_{r}$, and if these conditions are not met, datasets are either removed from the cluster or added based on their distance from the cluster centroid.
+These small adjustments are only made in cases where the clustering technique does not yield results that satisfy our requirements.
 IV Evaluation
@@
 Specifically, cosine similarity in combination with the $300$-dimensional vector embedding reduced the error rate in operator predictions, demonstrating that projecting datasets into this representation space and applying cosine similarity improves the prediction accuracy on the modelled operator.
-For the MLP classifier from sub-figures 5(c) and 5(d), the results illustrate that using vector embeddings of size $200$ and K-Means clustering produced the most accurate MLP classifier operators.
+For the MLP classifier (Figures 5(c) and 5(d)), the results illustrate that using vector embeddings of size $300$ and Cosine similarity- produced the most accurate MLP classifier operators.
 (a) SVM with SGD RMSE error loss
@@
 Apollo-SR $0.2$ SVM $68.70$ $22.81$ $2.10$ $1.16$
-$200$V Cosine MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
-$200$V SR-$0.4$ MLP $15.95$ $13.31$ $1.02$ $1.77$
+$300$V Cosine MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
+$300$V SR-$0.4$ MLP $15.95$ $13.31$ $1.02$ $1.77$
 Apollo-SR $0.1$ MLP $69.62$ $23.10$ $\mathbf{1.34}$ $1.14$

### v3

@@
-Data Analysis Prediction over Multiple Unseen Datasets: A Vector Embedding Approach
+Analytics Modelling over Multiple Datasets using Vector Embeddings
 Abstract
-The rapid growth in data volume and the increasing availability of datasets present analysts with the challenge of selecting high-quality data to optimize the performance of analytics operations.
-While choosing the best-quality data significantly improves accuracy and efficiency, it remains a complex task, particularly when dealing with a vast number of inputs.
-To tackle this challenge, we introduce a novel approach that predicts analytics outcomes by building a model based on datasets similar to the target dataset.
-Our method determines dataset similarity by mapping each dataset into a vector embedding representation.
-This transformation is achieved through our deep learning model, NumTabData2Vec, which processes entire datasets and projects them into a lower-dimensional vector space.
-Through experimental evaluation, we compare our framework’s prediction accuracy and execution time against a leading state-of-the-art modelling operator, demonstrating that our approach effectively predicts analytics outcomes.
-Additionally, our vectorization model can capture and differentiate diverse real-world scenarios by representing them in a compact vector space.
-Index Terms:
-Analytics Modelling, Vector embeddings, Vector Similarity, Data Quality
-I Introduction
+The massive increase in the data volume and dataset availability for analysts compels researchers to focus on data content and select high-quality datasets to enhance the performance of analytics operators.
+While selecting high-quality data significantly boosts analytical accuracy and efficiency, the exact process is very challenging given large-scale dataset availability.
+To address this issue, we propose a novel methodology that infers the outcome of analytics operators by creating a model from the available datasets.
+Each dataset is transformed to a vector embedding representation generated by our proposed deep learning model NumTabData2Vec, where similarity search are employed.
+Through experimental evaluation, we compare the prediction performance and the execution time of our framework to another state-of-the-art modelling operator framework, illustrating that our approach predicts analytics outcomes accurately, and increases speedup.
+Furthermore, our vectorization model can project different real-world scenarios to a lower vector embedding representation accurately and distinguish them.
+Keywords:
+Data Quality Analytics Modelling Vector embeddings Vector Similarity
+1 Introduction
 Big data technologies daily face the rapid evolution in volume as well as variety and velocity of processed data [12].
 Such big data characteristics routinely force analytics pipelines to underperform, requiring continuous maintenance and optimization.
-One major reason for this is bad quality of data [^1] https://tinyurl.com/de62sf48.
+One major reason for this is bad data quality [^1] https://tinyurl.com/de62sf48.
 Poor data quality leads to low data utilisation efficiency and even brings forth serious decision-making errors[6].
@@
 Data-centric Artificial Intelligence (AI) [30] emphasises on the quality, context, and structure of the data to improve its quality, as well as the analytical or machine learning (ML) algorithmic performance.
-Understating the data context properties such as data features, origins, relevance, and potential biases plays a critical role in modelling more accurate and reliable models.
+Understanding the data context properties, such as data features, origins, relevance, and potential biases, plays a critical role in modelling more accurate and reliable models.
 Data-centric AI prioritises the process of refining and enriching datasets to make them more suitable for real-world applications.
-In a similar way, a lot of scientists make the case that improving data quality by focusing on the content is a crucial factor in achieving better results [3].
-A plethora of available data sources and datasets in the organisations’ data centres and lakes pose a significant challenge: Deciding which datasets should be selected and passed into analytic workflows in order to produce more accurate results/predictions.
-While modern analytics workflows have evolved to incorporate diverse operators, the process of selecting the most suitable datasets using data-centric AI techniques to optimise a possibly complex computation is still under research [4].
-When dataset selection is left to human experts, it can result in low prediction performance and can be time-consuming.
-Equally costly can be the process of exhaustively executing over all (or many) available dataset inputs, in order to identify high-quality inputs according to business standards.
-In previous work [12], predicting the output of an analytics operator assuming a plethora of available input datasets was tackled via the creation of an all-pair similarity matrix, which, relative to the similarity function used, reflected the distance between datasets over a single characteristic (e.g., data distribution).
-To improve the big data analysis and utilize modern AI methods and systems, data or vector embeddings have been suggested.
-Data embedding vectorization [14, 15] aims at projecting data from a high-dimensional representation space into a more compact, lower dimensional space.
-By extracting meaningful patterns from the data features using deep learning, data is projected to a lower representation space.
-Using this space, one can find the most similar data points among a dataset.
-For the purpose of improving the accuracy of a modelled analytic operator (i.e., predict the outcome of a ML algorithm without actually executing it due to its cost), we propose a modelling framework that employs vector embeddings for dataset selection from a data lake store that contains a large number of available datasets.
-In order to predict an operator’s output for an “unseen” query dataset, our method considers only datasets that are qualitatively similar to the query, utilising a similarity search over the vector embeddings.
+Similarly, many researchers argue that prioritizing content-focused data quality is essential for achieving superior results [30].
+Yet, the plethora of available data sources and datasets in an organisation data repository poses a significant challenge: Deciding the most suitable datasets for analytics workflows to ensure accurate results/predictions.
+While modern analytics workflows incorporate diverse operators, optimising dataset selection using data-centric AI methods remains an active research area [15].
+When dataset selection is left to human experts, prediction performance drops, and it consumes more time.
+Equally costly and inefficient is the evaluation of all available datasets to identify high-quality inputs.
+In previous work [9], predicting the output of an analytics operator assuming a plethora of available input datasets was tackled via the creation of an all-pair similarity matrix, which, relative to the similarity function used, reflected the distance between datasets over a single data quality metric (e.g., data distribution).
+Data or vector embeddings have been proposed to enhance big data analysis and modern AI systems.
+Data embedding vectorization [22, 24] aims at projecting data from a high-dimensional representation space into a more compact, lower-dimensional space.
+Extracting meaningful information through data features using deep learning, data is projected to a lower representation space.
+To improve the accuracy of a modelled analytic operator (i.e., predict the outcome of a ML algorithm without actually executing it due to its cost), we propose a framework that uses vector embeddings for dataset selection from a large data lake repository.
+Our method predicts an operator’s output for an “unseen” query dataset, by selecting qualitatively similar datasets through similarity search over the vector embeddings.
 The selection of similar datasets reduces the prediction error, as well as the cost to model the operator, under the assumption that realistic analytical operators perform similarly under similar inputs.
-Similarity search is thus executed on the dataset vector embedding representations in a low-dimensional space.
-Each vector embedding is generated using our proposed deep learning method, NumTabData2Vec, which takes an entire dataset as input and transforms it into a vector embedding representation space.
-While previous methods split the dataset into chunks and produced a vector for each chunk or used the dataset metadata information, we use the entire dataset utilising all data record values.
-Our framework thus models an operator using similar datasets to the queried one.
-An interesting feature of our methodology is the ability of the NumTabData2Vec to efficiently distinguish with detail between different datasets and record properties in the lower vector representation space.
-Furthermore, the same vector representation is used for the modelling of multiple analytics operators, offering enhanced flexibility and time-efficiency.
+The embeddings are generated using our deep learning method, NumTabData2Vec, which processes entire tabular datasets rather than chunks or metadata, enabling efficient distinction between datasets and flexible modelling of multiple operators.
 Compared to similar previous work [9, 4], our work uses state-of-the-art data representation (vector embeddings) which are able to capture multiple data properties that can be used in order to assess similarity, namely record order, dataset size, data distribution, etc.
 The main contributions of our work can be summarised as follows:
-  • We introduce a framework for operator modelling in order to predict its outcome on an unseen tabular input dataset from a plethora of available ones.
+  • We introduce a framework for operator modelling (open-source prototype [^2] GitHub Repository) in order to predict its outcome on an unseen tabular input dataset from a plethora of available ones.
 Our method uses dataset vector embedding representations to improve the prediction performance via selecting the most relevant datasets to base its prediction upon.
   • We develop a deep learning model architecture that transforms an entire tabular dataset of numerical values to a vector embedding representation.
-  • We provide an experimental evaluation of our proposed methodology using multiple real-world scenarios and compare it directly to system Apollo [12, 13].
-Our evaluation illustrates that our proposed methodology can model operators with low prediction error by adaptively selecting similar quality datasets for each model.
-Moreover, it is able to do so with considerable amortized speed-ups.
-The evaluation of NumTabData2Vec indicates that it accurately projects any available dataset into a vector embedding representation and accurately models different dataset properties within the representation space.
-II Related Work
-Prior efforts largely concentrate on boosting/scaling an algorithm’s performance utilizing only the size of data input (record number) without examining its quality.
-As such, we also discuss works that try to identify the most suitable data features for the optimisation of analytic task operators.
+  • We provide an experimental evaluation of our proposed methodology using multiple real-world scenarios and compare it directly to the Apollo system [9, 4].
+Our evaluation illustrates that our methodology produces low prediction error by adaptively selecting similar quality datasets, achieving significant amortized speed-ups.
+NumTabData2Vec evaluation shows that it effectively projects datasets into vector embeddings while accurately capturing diverse dataset properties within the representation space.
+2 Related Work
+Prior efforts have focused on boosting algorithm performance by increasing data input (record number) rather than assessing quality.
+Consequently, we review works that identify optimal data features for analytic operator optimization.
 Vectorising data to lower embedding representation is a modern method that helps in identifying significant features across data types and datasets.
 As vector embeddings extract important features from data, we discuss studies that used the feature representation of data tuples to improve ML model prediction.
-II-A Data Quality
+2.1 Data Quality
 Big data applications aim to improve data quality by addressing various challenges.
-Dagger [8] enhances data quality by detecting errors in data pipelines using an SQL-like language and allowing code-handling for variables and arrays.
-ReClean [9] automates tabular data cleaning via reinforcement learning, optimizing tool selection for error repair.
+Dagger [26] enhances data quality by detecting pipeline errors using an SQL-like language, while ReClean [3] automates tabular data cleaning via reinforcement learning.
 IterClean [25] employs a large language model (LLM) to iteratively clean data by labelling initial tuples and using error detection, verification, and repair.
 In [8], data tuple quality is measured using Shapley values from game theory, with Truncated Monte Carlo Shapley and Gradient Shapley methods estimating a tuple’s value to a learning algorithm.
-Apollo [12, 13] is a content-based method that predicts the outcome of an analytic operator by leveraging dataset similarity.
-It involves three steps: creating a similarity matrix, projecting datasets to a lower-dimensional space, and modelling the operator using a small random subset of datasets.
+Apollo [9, 4] is a content-based method predicts analytic operator outcomes by leveraging dataset similarity through three steps: creating a similarity matrix, projecting datasets to a lower-dimensional space, and modelling the operator using a small random subset of datasets.
 Unlike Apollo, our approach selects the most relevant, high-quality datasets to model analytic operators, aiming to improve prediction performance.
-Moreover, modelling and similarity occur over the vector embeddings which are able to incorporate all dataset properties.
-In contrast, similarity functions used in [12] can only target a single property.
-II-B Dataset Selection Inference
+Additionally, our vector embeddings incorporate all dataset properties, whereas Apollo’s [9] similarity functions target only a single property.
+2.2 Dataset Selection Inference
 SOALA [11] selects optimal data features through online pairwise comparisons to maintain ML models over time.
@@
 Similarly, our framework uses dataset vector embeddings to select the most suitable datasets for modelling analytic operators or ML models, enhancing prediction accuracy.
-II-C Data Vectorization and Embeddings
+2.3 Data Vectorization and Embeddings
 The goal of data vectorization is to project high-dimensional data into a lower-dimensional vector space.
-Word2Vec [14] generates word embeddings using two neural network architectures, Continuous Bag of Words and Skip-Gram, leveraging word context.
-Graph2Vec [15] creates graph embeddings by dividing graphs into sub-graphs with a skip-graph model and aggregating their embeddings.
+Word2Vec [22] (using Continuous Bag of Words and Skip-Gram) [22] leverages word context to generate embeddings.
+Graph2Vec [24] creates graph embeddings by dividing graphs into sub-graphs with a skipgram model and aggregating their embeddings.
 ImageDataset2Vec [7] extracts meta-features from image datasets to generate embeddings, helping to select the most suitable classification algorithm.
@@
 Table2Vec [31] generates table embeddings by incorporating data features, metadata, and structural elements like captions and column headings.
+Mix2Vec [32], is an unsupervised deep neural network that projects mixed data into vector embeddings.
+In a clustering experiment like in their work on the common Adult dataset, our model outperformed Mix2Vec (recent method without publicly available code) by nearly $10\%$, demonstrating superior performance.
 Inspired by these methods, we designed a model that generates vector representations of tabular datasets over their record data values, not their metadata.
 Vector embeddings, which capture valuable information from data tuples, are widely used in classification tasks.
-TransTab [20] transforms data tuples into vector embeddings, applying transformer layers to encode features and predict classes, enhanced by supervised and self-supervised pretraining.
-FT-Transformer [21] and Res-Net architectures similarly use embeddings of categorical and numerical features, processed through transformer layers for class prediction.
-Tab-Transformer [22] embeds categorical features into vectors, normalizes continuous features, and combines them in a Multi-Layer Perceptron (MLP) to predict classes.
-Unlike these methods, which focus on tuple-level embeddings, our framework uses dataset-level vector embeddings to identify the most relevant datasets, improving analytic operator performance.
-III Methodology
+TransTab [28] encodes features with transformer layers to predict classes, leveraging supervised and self-supervised pretraining.
+FT-Transformer [10] and Res-Net architectures similarly use embeddings of categorical and numerical features, processed through transformer layers for class prediction, while Tab-Transformer [14] combines embedded categorical and normalized continuous features in an MLP for class prediction.
+Unlike these tuple-level approaches, our framework uses dataset-level embeddings to identify relevant datasets, enhancing analytic operator performance.
+3 Methodology
 Figure 1: Pipeline framework architecture
@@
 We also describe our approach for vectorizing tabular datasets, NumTabData2Vec.
-III-A Framework Architecture
-Consider a data lake store that contains a (possibly large) number $n$ of datasets $D=\left(d_{1},\ d_{2},\ d_{3},\ \ldots,\ d_{n}\right)$.
-Also, let us consider an analytics operator (e.g., a ML algorithm) $\Phi$ and an “unseen” dataset $D_{o}$.
-We assume that each $D_{i},1\leq i\leq n$ as well as $D_{o}$ consist of records with the same number of columns with numerical values only.
-Each dataset can, naturally, have a different number of records.
+3.1 Framework Architecture
+Consider a data lake repository that contains a (possibly large) number $n$ of structured tabular datasets $D=\left(d_{1},\ d_{2},\ d_{3},\ \ldots,\ d_{n}\right)$.
+Also, let us consider an analytics operator (e.g., a ML algorithm) $\Phi$ and an “unseen” dataset $D_{o}$ (from the same domain).
+We assume that each $D_{i},1\leq i\leq n$ as well as $D_{o}$ consist of records with numerical values only.
+Each dataset can, naturally, consist of different number of records.
 Operator $\Phi$ consumes a single such dataset as input to produce a single numerical output: $\Phi:D_{i}\rightarrow\mathbb{R}$.
-Our goal is to predict $\Phi(D_{o}),\forall D_{o}$ with a minimal cost and prediction error.
-Our premise is that the output of an analytical operator on any unseen $D_{o}$ can be accurately modelled and runtime efficiency improved by selecting a small subset of $D$.
-This can be achieved by identifying the most similar datasets to $D$, denoted: $D_{o}$: $D_{r}\subseteq D_{o}$.
+Our goal is to predict $\Phi(D_{o})$ with minimal cost and error by modelling the operator’s output for $D_{o}$ using a small subset of similar datasets $D_{r}\subseteq D$.
 Datasets in $D_{r}$ closely match $D_{o}$ in their properties (e.g., order, distribution, and size to name a few).
@@
 Figure 1 depicts the pipeline of our proposed framework.
-Datasets in $D$ are transformed into $k$-dimensional vectors via our NumTabData2Vec embedding scheme.
-Dataset vectorization needs no recomputation as the vectors are saved and can be reused for the rest of the framework steps.
-Each time a new dataset $D_{o}$ needs to be inferred relative to an analytics operator $\Phi$, its vector embedding is equally created.
-The datasets used for the creation of the model are selected via similarity search.
-Using different similarity functions, a small subset of datasets similar to $D_{o}$ are identified and used in order to create a model that predicts $\Phi(D_{o})$.
+Datasets in $D$ are transformed into $k$-dimensional vectors using our NumTabData2Vec scheme, and these embeddings are stored for reuse.
+Each time a $D_{o}$ needs to be inferred relative to an analytics operator $\Phi$, its vector embedding is created.
+The datasets used for the creation of the model are selected via similarity search to produce a small subset of relevant datasets.
+These chosen datasets are then used to model and predict $\Phi(D_{o})$, ensuring that only high-quality, pertinent data is processed.
 With this approach, our framework is utilizing “right quality” data in its inference mechanism, with irrelevant and extraneous datasets being excluded from the modelling process.
-The datasets are embedded by our NumTabData2Vec method.
-It takes an entire dataset as input and transforms it into a $k$-dimensional vector embedding representation space named $z$.
-The Vector embedding $z$, is a lower dimension representation of the dataset with the entire characteristics of the dataset being represented.
-Each dataset from $D$ is projected into such a $k$-dimensional vector embedding representation.
-The datasets are consequently represented by $D_{z}$ which contains all vector representations $\left(z_{1},\ z_{2},\ z_{3},\ \ldots,\ z_{n}\right)$.
-Dataset $D_{o}$ is also vectorized into $z_{o}$.
+The datasets are embedded by our NumTabData2Vec method, which transforms each entire dataset in $D$ into a k-dimensional vector $z$ that captures all its characteristics.
+Our framework operates seamlessly across diverse real-world scenarios without modification, requiring only the specification of a repository containing distinct numerical tabular datasets.
+The Vector embedding $z$ is a lower-dimension representation of the dataset with the entire characteristics of the dataset being encoded.
+Dataset $D_{o}$ is similarly embedded as $z_{o}$.
 Using the embedding representation $z_{o}$ and applying different similarity functions over the vector representations $D_{z}$, we may choose the most similar subset of $D$.
 The final step of the pipeline involves the operator modelling with any relevant method (e.g., Linear Regression, SVM, Multi-Layer Perceptron, etc.).
-This model can be used in order to infer the value of $\Phi(D_{o})$.
-III-B NumTabData2Vec
+This model is then used in order to infer the value of $\Phi(D_{o})$.
 Figure 2: NumTabData2Vec deep learning model architecture
-This method transforms a dataset $D_{i}$ to a lower dimension representation $z$ by using only the numerical values of the dataset’s records and excluding metadata, column names, and filenames.
-We present a deep learning model architecture based on the variational autoencoder (VAE) [5] concept.
-For each dataset, data is projected from a high dimensional space to a lower vector embedding representation space $z$, with $1\times k$ dimension.
+3.2 NumTabData2Vec
+This method transforms a dataset $D_{i}$ into a lower-dimensional vector $z$ using only its numerical values while excluding metadata like column names and filenames.
+We present a deep learning model architecture based on the variational autoencoder (VAE) [17] concept, that projects high-dimensional data into a $1\times k$ vector embedding dimension.
 The proposed method is thus defined as:
-$NumTabData2Vec\left(D_{i}\right)\rightarrow z,$ (1)
+$NumTabData2Vec\left(D_{i}\right)\rightarrow z,\$ (1)
 where the model takes an $m\times n$ dimensional numerical dataset and projects it to a lower $k$-dimension space $z$ ($k>1$).
-We wish for our method to be generally applicable and able to use any scenario dataset that learns to project a vector embedding representation during the training phase.
-Also, we wish this method to be able to learn this vector embedding representation using a diverse amount of data during the training phase, and execute thereafter without the need of further training, fine-tuning, or online-training methods.
-Finally, we need our scheme to be able to quickly and precisely extract the vector for every dataset input as well as deal with varying dataset dimensions without requiring any modifications.
-The deep learning model architecture is depicted in Figure 2, where a dataset $D_{i}$ with dimensions $m\times n$ passes through the encoder and is projected into a vector embedding representation $z$.
-The dataset has to be recreated since the decoder uses the same layers as the encoder but in the opposite direction.
-The learning of the vector representation $z$ is using the probabilistic encoder $q_{\phi}\left(z|x\right)$ distribution and the probabilistic decoder $p_{\theta}\left(x|z\right)$ distribution utilising the Kullback–Leibler (KL) divergence [6].
+We desire our method to be generally applicable to any dataset by learning to project a vector embedding during training.
+We also expect this method to learn vector embeddings from diverse data and operate without additional training or fine-tuning.
+Finally, we ought our scheme to be able to quickly and precisely extract vectors from every input while handling varying dataset dimensions without modifications.
+The deep learning model architecture is depicted in Figure 2, where a dataset $D_{i}$ with dimensions $m\times n$ passes through the encoder and is projected into a vector embedding representation $z$, then reconstructed by a decoder that mirrors the encode.
+The vector representation $z$ is learned using a probabilistic encoder $q_{\phi}\left(z|x\right)$, and decoder $p_{\theta}\left(x|z\right)$ that learns the distribution utilising the Kullback–Leibler (KL) divergence [18].
 To achieve that, the following condition:
-$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)$ (2)
+$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)\$ (2)
 of Kullback–Leibler (KL) divergence must be minimised.
-To learn the new vector representation $z$, the dataset must be reconstructed back from $z$ to their input format using the decoder, to verify if the vector representation is on a lower dimension in terms of the input dataset.
-This is the reconstruction loss.
-The loss function can be described as:
+To learn the new vector representation $z$, the dataset must be reconstructed back from $z$ to its input format using the decoder, to verify that the vector representation is compact.
+This reconstruction loss is part of the overall loss function, defined as:
 $\mathcal{L}_{\theta,\phi}\left(x\right)=\mathbb{E}_{q_{\phi}\left(z|x\right)}\left(\log{\left(p_{\theta}\left(x|z\right)\right)}\right)-D_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)$
 This loss function is called evidences lower bound (ELBO). While the KL divergence is minimised to learn the vector embedding representation $z$, the ELBO is maximized so the condition,
-$argmax\mathcal{L}_{\theta,\phi}\left(x\right)$ (4)
+$argmax\mathcal{L}_{\theta,\phi}\left(x\right)\$ (4)
 must be satisfied. The Decoder $p_{\theta}\left(x|z\right)$ is only used during the training phase to teach the encoder how to project the vector embedding representation $z$.
-The decoder takes as an input the dataset $D_{i}$, first extracts the feature embeddings from the input layer and passes them into the Transformer layers.
+The decoder extracts feature embeddings from the input dataset $D_{i}$ and processes them through Transformer layers.
 For the transformer layer we are using the pre-LN Transformer layer [29] instead of the traditional post-LN Transformer layer where the normalisation layer is employed inside the residual connection and before the prediction of the feed-forward layer.
-Our model design takes into account and uses the Pre-LN Transformer layer as it demonstrates improved learning outcomes with shorter training times [7].
-Following that, the transformed embedding space is projected onto a probabilistic distribution vector space $z$ using the mean ($\mu$) and standard deviation ($\sigma$).
-Then the transformed embedding space using $\mu$ and $\sigma$ are projected into a probabilistic distribution vector space $z$.
-The vector embedding space $z$ is a lower-dimensional representation of the dataset $D_{i}$ that preserves all the essential information needed to describe the dataset within the space $z$.
-The higher the $k$-dimension of the $z$ vector embedding representation, the higher and more accurate is this representation, as more features are extracted from the dataset [18].
-III-C Dataset Selection
+Following that, the transformed embedding space is projected into a probabilistic vector space z using the mean ($\mu$) and standard deviation ($\sigma$).
+This lower-dimensional space retains all essential information about $D_{i}$, and a higher dimension $k$ in $z$ leads to a more accurate representation by capturing additional features [16].
+3.3 Dataset Selection
 Algorithm 1 K-Means Clustering Algorithm for dataset selection
@@
 Different similarity functions are easily plugged into our pipeline.
-The first method uses the cosine similarity to determine how similar two datasets are in their vector embedding space $z$.
-Cosine similarity calculates the angle between the two vectors and a measure of similarity that is independent of the vector magnitudes.
-The higher its value, the most similar are the two datasets.
-The following equation describes the cosine similarity between vectors $A$ and $B$:
-$\cos\left(\theta\right)=\dfrac{A\cdot B}{\left\|A\right\|_{2}\left\|B\right\|_{2}}=\dfrac{\sum_{i=1}^{n}A_{i}B_{i}}{\sqrt{\sum_{i=1}^{n}A^{2}_{i}}\sqrt{\sum_{i=1}^{n}B^{2}_{i}}}$ (5)
-The alternative method involves calculating the Euclidean distance between the two vectors in order to determine how similar the two datasets are.
-Euclidean distance is particularly useful for capturing the geometric closeness between vectors and finding the most relevant datasets based on the organisation’s requirements.
-The Euclidean distance between vectors $\mathbf{u},\mathbf{v}$ is defined as:
-$d(\mathbf{u},\mathbf{v})=\sqrt{\sum_{i=1}^{n}(u_{i}-v_{i})^{2}}$ (6)
-The smaller the distance value the most similar are the datasets.
-Dataset selection using cosine similarity or euclidean distance selects a percentage of $\lambda,\lambda>0$ of the closest datasets to dataset $D_{o}$.
+The first method uses cosine similarity, which measures the angle between two vectors in the embedding space $z$ independent of their magnitudes, with a higher value indicating greater similarity.
+The alternative method calculates the Euclidean distance between two vectors to capture their geometric closeness and determine dataset similarity.
+This approach aids in selecting the most relevant datasets according to organizational requirements.
+The smaller the distance value then more similar are the datasets.
+Dataset selection using cosine similarity or euclidean distance selects a fraction of $\lambda,\lambda>0$ of the closest datasets to dataset $D_{o}$.
 The third approach involves utilising the K-Means [19, 21] clustering technique to choose the most relevant datasets.
 The datasets from $D$ are divided into $s$ ($s>1$) separate clusters, where datasets with similar features are grouped together in the same cluster based on similarity equations.
-The silhouette scores [25] are used to determine the number of clusters $s$ for each possible value of $S$.
+We determine the optimal number of clusters using silhouette scores [27].
 This is done by the following equation:
-$SilhouetteScore(v_{i})=\frac{b(v_{i})-a(v_{i})}{\max(a(v_{i}),b(v_{i}))}$ (7)
-where for each vector $v_{i}$ finds the mean intra-cluster distance ($a(v_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(v_{i})$) is the minimum average distance with the other vectors in the different cluster.
-Silhouette Score ranges from $-1$ to $1$ and the higher value defines the best $s$ number for clusters.
-Algorithm 1, shown how k-means clustering is created and how the most relevant datasets $D_{r}$ are selected based on the target dataset $D_{o}$.
-Using the most optimal $s$ value with the highest Silhouette score, K-means clustering is performed on the vector representations $V$ of each dataset, grouping datasets with same features into the same cluster.
-Next, the algorithm uses the vector $v_{o}$ of dataset $D_{o}$ to find the closest cluster centroid.
-All datasets in that cluster are defined as the relevant datasets $D_{r}$.
-These datasets are then used to model the analytics operator at hand.
+$SilhouetteScore(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max(a(z_{i}),b(z_{i}))},\$ (5)
+where for each vector $z_{i}$ computes the mean intra-cluster distance ($a(z_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(z_{i})$) is the minimum average distance with the other vectors in a different cluster.
+Silhouette Score ranges from $-1$ to $1$, and the higher value defines the best $s$ number for clusters.
+Algorithm 1, outlines the K-means process for selecting relevant datasets $D_{r}$ based on the target dataset $D_{o}$.
+Using the optimal s (with the highest Silhouette score), the vector representations $z$ of each dataset are clustered.
+Next, the algorithm uses the vector $z_{o}$ of dataset $D_{o}$ to find the closest cluster centroid.
+All datasets in that cluster are defined as the relevant datasets $D_{r}$, which are then used to model the analytics operator.
 However, we defined a maximum and minimum size for $D_{r}$, and if these conditions are not met, datasets are either removed from the cluster or added based on their distance from the cluster centroid.
 These small adjustments are only made in cases where the clustering technique does not yield results that satisfy our requirements.
-IV Evaluation
+4 Evaluation
 We compared our framework with Apollo [9, 4], which models analytic operators using data content.
@@
 RMSE is sensitive to outliers, while MAE is not; conversely, RMSE accounts for error direction, which MAE cannot.
-Speedup metrics are also used to evaluate how efficiently our framework models operator $\Phi$.
-Specifically, Speedup and Amortized Speedup measure the time required to approximate each operator versus exhaustively executing them on all datasets.
-Speedup is defined as $\frac{T{{}^{(i)}_{op}}}{T{{}^{(i)}_{SimOp}+T_{vec}+T_{sim}+T_{pred}}}$, where $T{{}^{(i)}_{op}}$ is the time to execute operator $i$ on all datasets, $T{{}^{(i)}_{SimOp}}$ is the time to model the operator with datasets from similarity search, $T_{vec}$ is the vector embedding computation time, $T_{sim}$ is the similarity search time, and $T_{pred}$ is the prediction time for $D_{o}$.
-Amortized speedup includes dataset vectorization, performed once per data lake for multiple operators (in our case two operators).
-We also evaluate our dataset vectorization model, NumTabData2Vec, which projects datasets into vector embedding space $z$.
-Three versions were built with vector sizes of $100$, $200$, and $300$, each featuring eight transformer layers.
-The models were trained for 50 epochs on four NVIDIA A10 GPUs in parallel.
-IV-A Evaluation Setup
+We further assess efficiency using Speedup and Amortized Speedup metrics, where Speedup is defined as $\frac{T{{}^{(i)}_{op}}}{T{{}^{(i)}_{SimOp}+T_{vec}+T_{sim}+T_{pred}}}$, where $T{{}^{(i)}_{op}}$ is the time to execute operator $i$ on all datasets, $T{{}^{(i)}_{SimOp}}$ is the time to model the operator with datasets from similarity search, $T_{vec}$ is the vector embedding computation time, $T_{sim}$ is the similarity search time, and $T_{pred}$ is the prediction time for $D_{o}$.
+Amortized speedup including one-time vectorization per data lake across multiple operators.
+Three variants with vector sizes 100, 200, and 300 (each with eight transformer layers) were trained for $100$ epochs on four NVIDIA A10101010 GPUs.
+More experimental evaluation results can be found in the extended version of this work [20].
+4.1 Evaluation Setup
 Our framework is deployed over an AWS EC2 virtual machine server running with 48 vCPUs of AMD EPYC 7R32 processors at 2.40GHz, and four A10s GPUs with 24GB of memory each, $192GB$ of RAM memory, and $2TB$ of storage, running over Ubuntu 24.4 LTS.
 Our code is written in Python (v.3.9.1) and PyTorch modules (v.2.4.0).
-Apollo was deployed in a virtual machine with 8 VCPUs Intel Xeon E5-2630 @ 2.30GHz, $64GB$ of RAM memory, and $250GB$ of storage, running Ubuntu 24.4 LTS like in their experimental evaluation.
-IV-B Datasets
-TABLE I: Dataset properties for experimental evaluation
+Apollo was deployed on the same AWS EC2 virtual machine server, utilizing only the vCPUs and RAM, as it does not require a GPU for execution.
+4.2 Datasets
+Table 1: Dataset properties for experimental evaluation
 Dataset Name # Files # Tuples # Columns
@@
 Weather [2] $49$ $516$ 7
-We evaluated our framework using four diverse datasets to represent real-world scenarios.
-Table I summarizes these datasets and their attributes.
-The vectorization module, NumTabData2Vec, was trained on data separate from the experimental evaluation data, split $60\%$ for training and $40\%$ for testing.
-The Household Power Consumption (HPC) dataset [28] contains electric power usage measurements from a household in Sceaux, France.
-It includes $401$ datasets, each with $2051$ tuples and seven features recorded at one-minute intervals.
-The Adult dataset [29], commonly used for binary classification, predicts whether an individual earns more or less than $50K$ annually.
-It comprises $100$ datasets, each with $228$ individuals and various socio-economic features.
-The Stock Market dataset [30] includes daily NASDAQ stock prices obtained from Yahoo Finance, with $508$ datasets.
-Each dataset contains $13$ to $1959$ tuples, each describing seven feature attributes.
-The Weather dataset [31] provides hourly weather measurements from $36$ U.S.
-cities between $2012$ and $2017$, split into $49$ datasets, each with $516$ tuples and seven features.
-(a) Linear Regression RMSE error loss
-(b) Linear Regression MAE error loss
-(c) MLP for Regression RMSE error loss
-(d) MLP for Regression MAE error loss
+We evaluated our framework using four real-world datasets (see Table 1).
+The NumTabData2Vec module was trained on separate data ($60\%$ training, $40\%$ testing).
+The Household Power Consumption (HPC) dataset [13] contains $401$ datasets with $2051$ tuples and seven features recorded at one-minute intervals of electric power usage measurements.
+The Adult dataset [5], used for binary classification, predicts income levels and includes $100$ datasets with $228$ individuals and socio-economic features.
+The Stock Market dataset [1] consists of $508$ datasets with $13$ to $1959$ tuples describing daily NASDAQ stock prices.
+Weather dataset [2] provides hourly measurements from $36$ U.S.
+cities (2012–2017), split into $49$ datasets with $516$ tuples and seven features.
+Any categorical feature column in all datasets is transformed to numerical data by one-hot encoding.
+These datasets were selected to demonstrate our framework’s ability to perform consistently across diverse real-world scenarios.
+Our framework was evaluated by predicting the outputs of various ML operators without directly executing them.
+Datasets were projected into $k$-dimensional spaces with vector dimensions of $100$, $200$, and $300$.
+For regression datasets (Household Power Consumption and Stock Market), we modelled Linear Regression (LR) and Multi-Layer Perceptron (MLP), while for classification datasets (Weather and Adult), we modelled Support Vector Machine (SVM) and MLP classifiers.
+Each experiment has executed 10 times, and we report the average error loss and speedup.
+4.3 Evaluation Results
+Figures 4, 4, 6, and 6 present the evaluation results of different similarity search methods across vector embedding spaces of sizes $100$, $200$, and $300$ (green, blue, and grey bars, respectively).
+In each sub-figure, the y-axis represents the error loss value, while the x-axis displays the similarity search method applied over the vector embeddings.
+Figures 4 and 4 display results for the Stock Market and Household Power Consumption datasets, with MLP regression in the bottom sub-figure and LR in the top.
+Figures 6 and 6 show results for the Weather and Adult datasets, with SVM (SGD) in the top sub-figure and MLP classifier in the bottom.
+Left sub-figures use RMSE loss, while right sub-figures use MAE loss.
 Figure 3: Household power consumption dataset prediction error loss
-Our framework was evaluated by registering the accuracy of predicting the output of various ML operators over multiple datasets in $D$ without actually executing the operator on them.
-To evaluate our scheme and its parameters, we use all four datasets, ranging the size of the produced vectors as well as the similarity functions used.
-We project all datasets into $k$-dimensional spaces with varying vector dimensions ($100$, $200$, and $300$).
-For each dataset in Table I, we model different operators: For the regression datasets (Household Power Consumption and Stock Market), we model Linear Regression (LR) and Multi-Layer Perceptron (MLP) operators; for the classification datasets (Weather and Adult), we model the Support Vector Machine (SVM) and MLP classifier operators.
-Each experiment has been executed $10$ times and we report the average of the error loss, as well as the speedup.
-(a) Linear Regression RMSE error loss
-(b) Linear Regression MAE error loss
-(c) MLP for Regression RMSE error loss
-(d) MLP for Regression MAE error loss
 Figure 4: Stock market dataset prediction error loss
-IV-C Evaluation Results
-Figures 3, 4, 5, and 6 present the evaluation results for each method, comparing the performance of different similarity search techniques across various vector embedding representation spaces.
-The red (with hatches), brown, and blue bars correspond to vector embeddings of size 100, 200, and 300 respectively.
-In each sub-figure, the y-axis represents the error loss value, while the x-axis displays the similarity search method applied over the vector embeddings.
-Figures 3 and 4 show the results for the Stock market and Household power consumption datasets, where the bottom sub-figure demonstrates the MLP regression model, and the top sub-figure presents the LR model.
-Figures 5 and 6 depict the evaluation results for the Weather and Adult datasets.
-In these Figures, the top sub-figure shows the SVM with SGD results, while the bottom sub-figure shows the MLP classifier.
-The left sub-figures in all Figures use the RMSE loss function, whereas the right sub-figures use the MAE loss function.
-Figure 3, we show, for the HPC dataset, shows as increase the vector dimension size there is slightly lower prediction error for all the operator modelling.
-While for different similarity methods did not result in any significant differences in the prediction error loss for all the operator modelling.
+Figure 4, for the HPC dataset, shows as increase the vector dimension size there is slightly lower prediction error for all the operator modelling.
+Different similarity methods do not result in any significant differences in the prediction error loss for all the operator modelling.
 This suggests that, regardless the similarity selection method, our framework effectively selects the most optimal subset of data to improve model predictions on the unseen input dataset $D_{o}$.
@@
 However, due to the inherent volatility in Stock market data from different days, all models in the stock market dataset experiments exhibit high loss values.
-In the weather dataset, the SVM operator results from sub-figures 5(a) and 5(b) show that using $300$ vectors in the representation space consistently led to more accurate operator models across all similarity methods.
+In the weather dataset, the SVM operator results (Figures 5(a) and 5(b)) show that using $300$ sized vectors in the representation space consistently led to more accurate operator models across all similarity methods.
 Specifically, cosine similarity in combination with the $300$-dimensional vector embedding reduced the error rate in operator predictions, demonstrating that projecting datasets into this representation space and applying cosine similarity improves the prediction accuracy on the modelled operator.
 For the MLP classifier (Figures 5(c) and 5(d)), the results illustrate that using vector embeddings of size $300$ and Cosine similarity- produced the most accurate MLP classifier operators.
-(a) SVM with SGD RMSE error loss
-(b) SVM with SGD MAE error loss
-(c) MLP RMSE error loss
-(d) MLP MAE error loss
 Figure 5: Weather dataset prediction error loss
+Figure 6: Adult dataset prediction error loss
 On the other hand, the Adult dataset shows the lowest error rates, with error loss values consistently below $0.5$ across all vector embedding dimensions and similarity search methods (see Figure 6).
@@
 This confirms that a higher number of vector dimensions leads to more accurate predictions, consistent with findings in previous research [22].
-(a) SVM with SGD RMSE error loss
-(b) SVM with SGD MAE error loss
-(c) MLP RMSE error loss
-(d) MLP MAE error loss
-Figure 6: Adult dataset prediction error loss
+(a) Adult Dataset MLP Operator RMSE error loss
+(b) HPC dataset LR Operator RMSE error loss
+Figure 7: Sampling Ratio prediction results
 We conducted an experimental evaluation using the Sampling Ratio (SR) approach, similar to Apollo [9], but employed neural networks built from the vector embeddings of each dataset.
@@
 In each sub-figure the y-axis represents the RMSE prediction error loss while the x-axis denotes the vector dimension
-(a) Adult Dataset MLP Operator RMSE error loss
-(b) HPC dataset LR Operator RMSE error loss
-Figure 7: Sampling Ratio prediction results
 Both experiments demonstrate that as the vector embedding dimension increases, coupled with a larger sampling ratio (SR) value, there is a slight decrease in the prediction error loss.
@@
 Likewise, in the Adult dataset, the SR approach delivered the poorest performance, with nearly $90\%$ lower prediction accuracy compared to the similarity search methods.
-TABLE II: Evaluation results of our framework exported analytic operator with lowest prediction error in comparison with Apollo
+Table 2: Evaluation results of our framework exported analytic operator with lowest prediction error in comparison with Apollo
 Dataset Name Method Operator RMSE MAE Speedup Amortized Speedup
@@
 Apollo-SR $0.2$ MLP $70.16$ $25.74$ $1.05$ $1.31$
-Table II compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$.
+For the HPC and Weather datasets, the SR approach was approximately $15\%$ less accurate in operator prediction compared to all similarity search methods, even as vector embedding dimensions increased.
+In contrast, the Stock dataset exhibited a significantly larger discrepancy, with the SR approach performing about $70\%$ worse in prediction accuracy across all vector embedding dimensions.
+Similarly, in the Adult dataset, the SR approach recorded the poorest performance, delivering nearly $90\%$ worse prediction accuracy compared to the similarity search methods.
+Table 2 compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$.
 Methods $100$V, $200$V, and $300$V denote vector embedding dimensions.
 The lowest prediction errors align with our pipeline’s similarity search method.
-Apollo outperforms our framework on the Stock dataset for the LR analytic operator at SR equals with $0.1$ (with $50\%$ and $6\%$ improvements for RMSE and MAE, respectively).
+Apollo outperforms our framework on the Stock dataset for the LR analytic operator with the smallest amount of SR.
 However, our framework excels with the MLP regression operator, improving RMSE and MAE by $20\%$ and $17\%$, respectively.
@@
 However, the similarity search method outperforms both Apollo and the SR approach in prediction accuracy and amortized speedup, establishing its clear superiority across most datasets and operator scenarios.
-IV-D NumTabData2Vec Evaluation Results
+4.4 NumTabData2Vec Evaluation Results
+Table 3: Similarity between vectors of different datasets scenarios
+NumTabData2Vec Vector Size Similarity
+$100$ $0.54$
+$200$ $0.18$
+$300$ $0.16$
 Figure 8: Vector representation for each dataset from NumTabData2Vec
-TABLE III: Similarity between vectors of different datasets scenarios
-Model Name Similarity
-NumTabData2Vec $100$ Vector size $0.54$
-NumTabData2Vec $200$ Vector size $0.18$
-NumTabData2Vec $300$ Vector size $0.16$
-Our proposed model, NumTabData2Vec, was evaluated to determine its ability to distinguish dataset scenarios based on qualitative differences.
-For each scenario, $n$ random datasets were selected, and their vector embeddings averaged to represent the scenario.
-These embeddings, initially in a 200-dimensional space, were projected into 3D using PCA and are shown in Figure 8.
-The figure illustrates that each dataset scenario occupies a distinct space, with minimal overlap or clustering.
-This demonstrates that NumTabData2Vec effectively distinguishes datasets, outperforming prior methods like Word2Vec and Graph2Vec [14, 15], which achieved similar accuracy but on different data types (e.g., words, graphs) rather than entire datasets.
-Table III further highlights the average cosine similarity between dataset embeddings, showing greater dissimilarity as vector dimensions increase.
+Figure 9: Synthetic data vector embedding representation
+Our proposed model, NumTabData2Vec,was evaluated for its ability to distinguish dataset scenarios based on qualitative differences.
+For each scenario, $n$ random datasets were reduced to 3D vector embeddings using PCA, as shown in Figure 9, which demonstrates NumTabData2Vec’s ability to distinguish datasets with minimal overlap across contexts.
+Unlike prior methods [22, 24], which focus on text or graphs, NumTabData2Vec applies to entire datasets.
+Table 3 further highlights the average cosine similarity between dataset embeddings, showing greater dissimilarity as vector dimensions increase.
 However, results suggest that dimensions between $100$ and $300$ are sufficient for accurate distinction, avoiding the need for larger vector sizes.
-Figure 9: Synthetic data vector embedding representation
+Table 4: NumTabData2Vec execution time for different dataset dimensions and different vector sizes
+# of columns # of rows $50$ Vectors Execution time $100$ Vectors Execution time $200$ Vectors Execution time $3$ $100$ $0.0004$ sec $0.00042$ sec $0.00051$ sec $3$ $500$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1000$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1500$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $3$ $1800$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $10$ $100$ $0.0004$ sec $0.0004$ sec $0.00057$ sec $10$ $500$ $0.00039$ sec $0.0004$ sec $0.00051$ sec $10$ $1000$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $10$ $1500$ $0.00041$ sec $0.00042$ sec $0.00055$ sec $10$ $1800$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $20$ $100$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $500$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $1000$ $0.00042$ sec $0.00043$ sec $0.00052$ sec $20$ $1500$ $0.00043$ sec $0.00044$ sec $0.00054$ sec $20$ $1800$ $0.00044$ sec $0.00044$ sec $0.00054$ sec
+Figure 10: HPC Dataset vector embedding representation with addition of Noise
+Figure 11: HPC Dataset vector embedding representation with addition of Noise in the first column
 To evaluate NumTabData2Vec’s ability to distinguish datasets with varying row and column counts, we generated synthetic numerical tabular datasets of different dimensions and vectorized them.
@@
 These results indicate that our method effectively distinguishes datasets based on size during vectorization.
-TABLE IV: NumTabData2Vec execution time for different dataset dimensions and different vector sizes
-# of columns # of rows $50$ Vectors Execution time $100$ Vectors Execution time $200$ Vectors Execution time $3$ $100$ $0.0004$ sec $0.00042$ sec $0.00051$ sec $3$ $500$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1000$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1500$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $3$ $1800$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $10$ $100$ $0.0004$ sec $0.0004$ sec $0.00057$ sec $10$ $500$ $0.00039$ sec $0.0004$ sec $0.00051$ sec $10$ $1000$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $10$ $1500$ $0.00041$ sec $0.00042$ sec $0.00055$ sec $10$ $1800$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $20$ $100$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $500$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $1000$ $0.00042$ sec $0.00043$ sec $0.00052$ sec $20$ $1500$ $0.00043$ sec $0.00044$ sec $0.00054$ sec $20$ $1800$ $0.00044$ sec $0.00044$ sec $0.00054$ sec
+To evaluate NumTabData2Vec’s ability to distinguish datasets by distribution and order, Gaussian noise was added to $l\%$ of tuples in an HPC dataset.
+Figure 11 shows the original and noise-modified datasets projected to 2D using PCA, with greater noise causing larger shifts in the representation space.
+This demonstrates the model’s effectiveness in capturing distribution differences and distinguishing datasets based on ordering.
+To assess fine-grained distinctions, we repeated the experiment by adding Gaussian noise exclusively to the first column of the dataset.
+Figure 11 shows the 2D vector space, where g in the bullet caption indicates the noise level.
+As noise increases, the representation shifts further from the original dataset, though it remains closer than in Figure 11, with points more tightly grouped in 2D space.
 To evaluate how dataset dimensions affect the execution time of NumTabData2Vec, we created synthetic datasets with varying numbers of rows ($100$, $500$, $1000$, $1500$, and $1800$) and columns ($3$, $10$, and $20$).
 These datasets were vectorized into different dimensions, and the execution times were recorded.
-Table IV shows that increasing the k-dimension requires approximately $20\%$ more time to generate the vector embeddings.
+Table 4 shows that increasing the k-dimension requires approximately $20\%$ more time to generate the vector embeddings.
 This is expected, as a higher k-dimension involves more hyperparameters, which naturally increases computation time.
@@
 This is because larger datasets require the extraction of more features, which has a modest impact on the model’s execution time.
-Figure 10: HPC Dataset vector embedding representation with addition of Noise
-To evaluate NumTabData2Vec’s ability to distinguish datasets based on different properties like distribution and order, we introduced Gaussian noise to random $l\%$ of data tuples in an HPC dataset.
-Figure 10 visualises the original and noise-modified datasets, projected from a 200-dimensional space to 2D using PCA.
-Each bullet caption g denotes the percentage of Gaussian noise added in the dataset.
-As noise increases, the representation space shifts further from the original dataset, indicating that NumTabData2Vec effectively captures distribution differences.
-Additionally, since the HPC dataset has an inherent order, the model’s sensitivity to noise demonstrates its ability to distinguish datasets based on ordering as well.
-Figure 11: HPC Dataset vector embedding representation with addition of Noise in the first column
-To evaluate how fine-grained as distinction can be, we introduced noise into a single column and repeated the previous experiment, with the difference being that noise was added exclusively to the first column.
-Figure 11 visualizes the dataset’s 2D vector space.
-The amount of Gaussian noise added to the dataset’s first column is indicated by g in the bullet caption.
-The results show that as more noise is introduced to the column, the vector representation moves further away from the original dataset.
-In contrast to the previous experiment shown in Figure 10, the noisy dataset’s representation stays closest to the original when only a single column is modified.
-Also in this experiment the dataset points in the 2-dimension are more grouped between them instead the previous experiment.
-V Limitations
+5 Limitations
 There exist a number of limitations in our work as we described it.
@@
 Finally, our proposed NumData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples.
-This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available).
-VI Conclusion
+This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available for a budget GPU).
+6 Conclusion
 In this paper, we presented a novel framework for the modelling of an analytic operator (such as a ML algorithm) when a large number of input data is available and thus no brute-force execution can be performed.
@@
 We also demonstrated that NumData2Vec can effectively detect when noise is introduced into a dataset.
-VII ACKNOWLEDGMENTS
+{credits}
+6.0.1 Acknowledgements
 This work is partially supported by the project RELAX-DN, funded by the European Union under Horizon Europe 2021-2027 Framework Programme Grant Agreement number 101072456.

### v4

@@
 The proposed method is thus defined as:
-$NumTabData2Vec\left(D_{i}\right)\rightarrow z,\$ (1)
+$NumTabData2Vec\left(D_{i}\right)\rightarrow z,\\$ (1)
 where the model takes an $m\times n$ dimensional numerical dataset and projects it to a lower $k$-dimension space $z$ ($k>1$).
@@
 To achieve that, the following condition:
-$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)\$ (2)
+$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)\\$ (2)
 of Kullback–Leibler (KL) divergence must be minimised.
@@
 This loss function is called evidences lower bound (ELBO). While the KL divergence is minimised to learn the vector embedding representation $z$, the ELBO is maximized so the condition,
-$argmax\mathcal{L}_{\theta,\phi}\left(x\right)\$ (4)
+$argmax\mathcal{L}_{\theta,\phi}\left(x\right)\\$ (4)
 must be satisfied. The Decoder $p_{\theta}\left(x|z\right)$ is only used during the training phase to teach the encoder how to project the vector embedding representation $z$.
@@
 7:   Assign $\mathbf{z}_{i}$ to the nearest centroid based on Euclidean distance:
-Assign $\text{Assign }\mathbf{z}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|^{2}$
+$\text{Assign }\mathbf{z}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|^{2}$
 8:  end for
@@
 This is done by the following equation:
-$SilhouetteScore(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max(a(z_{i}),b(z_{i}))},\$ (5)
+$SilhouetteScore(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max(a(z_{i}),b(z_{i}))},\\$ (5)
 where for each vector $z_{i}$ computes the mean intra-cluster distance ($a(z_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(z_{i})$) is the minimum average distance with the other vectors in a different cluster.
-Silhouette Score ranges from $-1$ to $1$, and the higher value defines the best $s$ number for clusters.
+Silhouette Score ranges from 1-1 to $1$, and the higher value defines the best $s$ number for clusters.
 Algorithm 1, outlines the K-means process for selecting relevant datasets $D_{r}$ based on the target dataset $D_{o}$.
@@
 Amortized speedup including one-time vectorization per data lake across multiple operators.
-Three variants with vector sizes 100, 200, and 300 (each with eight transformer layers) were trained for $100$ epochs on four NVIDIA A10101010 GPUs.
+Three variants with vector sizes 100, 200, and 300 (each with eight transformer layers) were trained for $100$ epochs on four NVIDIA A1010 GPUs.
 More experimental evaluation results can be found in the extended version of this work [20].
@@
 Table 1: Dataset properties for experimental evaluation
-Dataset Name # Files # Tuples # Columns
+Dataset Name
+# Files
+# Tuples
+# Columns
 Household Power
@@
 Adult [5] $100$ $228$ 14
-Stocks [1] $508$ $1959-13$ 7
+Stocks [1] $508$ 1959131959-13 7
 Weather [2] $49$ $516$ 7
@@
 Table 2: Evaluation results of our framework exported analytic operator with lowest prediction error in comparison with Apollo
-Dataset Name Method Operator RMSE MAE Speedup Amortized Speedup
-Household Power Consumption $300$V Cosine LR $\mathbf{6.61}$ $\mathbf{5.42}$ $0.0017$ $\mathbf{1.99}$
-$300$V SR-$0.2$ LR $7.77$ $6.66$ $0.0018$ $1.42$
-Apollo-SR $0.1$ LR $2968.01$ $2352.55$ $\mathbf{0.015}$ $0.024$
-Apollo-SR $0.2$ LR $2811.49$ $2229.50$ $0.015$ $0.024$
-$300$V K-Means MLP Regr. $\mathbf{6.70}$ $\mathbf{3.38}$ $0.9249$ $\mathbf{1.99}$
-Apollo-SR $0.1$ MLP Regr. $3322.05$ $2606.99$ $2.38$ $1.74$
-Apollo-SR $0.2$ MLP Regr. $3850.01$ $2609.36$ $\mathbf{2.38}$ $1.74$
-Stock $300$V Cosine LR $306382.28$ $125335.65$ $0.00085$ $\mathbf{1.91}$
-$300$V SR-$0.4$ LR $21861625.91$ $5674215.265$ $0.00087$ $0.33$
-Apollo-SR $0.1$ LR $\mathbf{153665.92}$ $\mathbf{118236.48}$ $\mathbf{0.00093}$ $0.00096$
-Apollo-SR $0.2$ LR $166844.95$ $133306.68$ $0.00093$ $0.00096$
-$300$V Cosine MLP Regr. $\mathbf{140236.47}$ $\mathbf{123571.12}$ $0.63$ $\mathbf{1.91}$
-Apollo-SR $0.1$ MLP Regr. $175150.82$ $145123.09$ $\mathbf{0.93}$ $0.96$
-Apollo-SR $0.2$ MLP Regr. $174390.81$ $146338.73$ $0.93$ $0.96$
-Weather $300$V Cosine SVM SGD $\mathbf{14.13}$ $\mathbf{7.63}$ $1.06$ $\mathbf{22.8}$
-Apollo-SR $0.1$ SVM $69.51$ $25.52$ $\mathbf{2.10}$ $1.16$
-Apollo-SR $0.2$ SVM $68.70$ $22.81$ $2.10$ $1.16$
-$300$V Cosine MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
-$300$V SR-$0.4$ MLP $15.95$ $13.31$ $1.02$ $1.77$
-Apollo-SR $0.1$ MLP $69.62$ $23.10$ $\mathbf{1.34}$ $1.14$
-Apollo-SR $0.2$ MLP $673.56$ $\mathbf{84.70}$ $1.32$ $1.14$
-Adult $300$V Cosine SVM SGD $\mathbf{0.36}$ $\mathbf{0.2}$ $0.37$ $\mathbf{2.78}$
-Apollo-SR $0.1$ SVM $68.32$ $22.95$ $\mathbf{0.75}$ $0.85$
-Apollo-SR $0.2$ SVM $68.88$ $22.88$ $0.74$ $0.85$
-$300$V K-Means MLP $\mathbf{0.36}$ $\mathbf{0.19}$ $0.30$ $2.78$
-$300$V SR-$0.2$ MLP $6.01$ $6.00$ $0.54$ $\mathbf{3.54}$
-Apollo-SR $0.1$ MLP $71.11$ $26.51$ $\mathbf{1.07}$ $1.31$
-Apollo-SR $0.2$ MLP $70.16$ $25.74$ $1.05$ $1.31$
+Dataset
+Name
+Method Operator RMSE MAE Speedup Amortized Speedup
+Household
+Power
+Consumption
+300300V Cosine
+LR $\mathbf{6.61}$ $\mathbf{5.42}$ $0.0017$ $\mathbf{1.99}$
+300300V SR-$0.2$
+LR $7.77$ $6.66$ $0.0018$ $1.42$
+Apollo-SR $0.1$
+LR $2968.01$ $2352.55$ $\mathbf{0.015}$ $0.024$
+Apollo-SR $0.2$
+LR $2811.49$ $2229.50$ $0.015$ $0.024$
+300300V K-Means
+MLP Regr. $\mathbf{6.70}$ $\mathbf{3.38}$ $0.9249$ $\mathbf{1.99}$
+Apollo-SR $0.1$
+MLP Regr. $3322.05$ $2606.99$ $2.38$ $1.74$
+Apollo-SR $0.2$
+MLP Regr. $3850.01$ $2609.36$ $\mathbf{2.38}$ $1.74$
+Stock
+300300V Cosine
+LR $306382.28$ $125335.65$ $0.00085$ $\mathbf{1.91}$
+300300V SR-$0.4$
+LR $21861625.91$ $5674215.265$ $0.00087$ $0.33$
+Apollo-SR $0.1$
+LR $\mathbf{153665.92}$ $\mathbf{118236.48}$ $\mathbf{0.00093}$ $0.00096$
+Apollo-SR $0.2$
+LR $166844.95$ $133306.68$ $0.00093$ $0.00096$
+300300V Cosine
+MLP Regr. $\mathbf{140236.47}$ $\mathbf{123571.12}$ $0.63$ $\mathbf{1.91}$
+Apollo-SR $0.1$
+MLP Regr. $175150.82$ $145123.09$ $\mathbf{0.93}$ $0.96$
+Apollo-SR $0.2$
+MLP Regr. $174390.81$ $146338.73$ $0.93$ $0.96$
+Weather
+300300V Cosine
+SVM SGD
+$\mathbf{14.13}$ $\mathbf{7.63}$ $1.06$ $\mathbf{22.8}$
+Apollo-SR $0.1$
+SVM $69.51$ $25.52$ $\mathbf{2.10}$ $1.16$
+Apollo-SR $0.2$
+SVM $68.70$ $22.81$ $2.10$ $1.16$
+300300V Cosine
+MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
+300300V SR-$0.4$
+MLP $15.95$ $13.31$ $1.02$ $1.77$
+Apollo-SR $0.1$
+MLP $69.62$ $23.10$ $\mathbf{1.34}$ $1.14$
+Apollo-SR $0.2$
+MLP $673.56$ $\mathbf{84.70}$ $1.32$ $1.14$
+Adult
+300300V Cosine
+SVM SGD
+$\mathbf{0.36}$ $\mathbf{0.2}$ $0.37$ $\mathbf{2.78}$
+Apollo-SR $0.1$
+SVM $68.32$ $22.95$ $\mathbf{0.75}$ $0.85$
+Apollo-SR $0.2$
+SVM $68.88$ $22.88$ $0.74$ $0.85$
+300300V K-Means
+MLP $\mathbf{0.36}$ $\mathbf{0.19}$ $0.30$ $2.78$
+300300V SR-$0.2$
+MLP $6.01$ $6.00$ $0.54$ $\mathbf{3.54}$
+Apollo-SR $0.1$
+MLP $71.11$ $26.51$ $\mathbf{1.07}$ $1.31$
+Apollo-SR $0.2$
+MLP $70.16$ $25.74$ $1.05$ $1.31$
 For the HPC and Weather datasets, the SR approach was approximately $15\%$ less accurate in operator prediction compared to all similarity search methods, even as vector embedding dimensions increased.
@@
 Table 2 compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$.
-Methods $100$V, $200$V, and $300$V denote vector embedding dimensions.
+Methods 100100V, 200200V, and 300300V denote vector embedding dimensions.
 The lowest prediction errors align with our pipeline’s similarity search method.
@@
 Table 3: Similarity between vectors of different datasets scenarios
-NumTabData2Vec Vector Size Similarity
-$100$ $0.54$
-$200$ $0.18$
-$300$ $0.16$
+NumTabData2Vec
+Vector Size
+Similarity
+$100$
+$0.54$
+$200$
+$0.18$
+$300$
+$0.16$
 Figure 8: Vector representation for each dataset from NumTabData2Vec
@@
 Table 4: NumTabData2Vec execution time for different dataset dimensions and different vector sizes
-# of columns # of rows $50$ Vectors Execution time $100$ Vectors Execution time $200$ Vectors Execution time $3$ $100$ $0.0004$ sec $0.00042$ sec $0.00051$ sec $3$ $500$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1000$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1500$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $3$ $1800$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $10$ $100$ $0.0004$ sec $0.0004$ sec $0.00057$ sec $10$ $500$ $0.00039$ sec $0.0004$ sec $0.00051$ sec $10$ $1000$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $10$ $1500$ $0.00041$ sec $0.00042$ sec $0.00055$ sec $10$ $1800$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $20$ $100$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $500$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $1000$ $0.00042$ sec $0.00043$ sec $0.00052$ sec $20$ $1500$ $0.00043$ sec $0.00044$ sec $0.00054$ sec $20$ $1800$ $0.00044$ sec $0.00044$ sec $0.00054$ sec
+# of columns # of rows $50$ Vectors Execution time $100$ Vectors Execution time $200$ Vectors Execution time 33 $100$ $0.0004$ sec $0.00042$ sec $0.00051$ sec 33 $500$ $0.0004$ sec $0.00041$ sec $0.00049$ sec 33 $1000$ $0.0004$ sec $0.00041$ sec $0.00049$ sec 33 $1500$ $0.0004$ sec $0.00041$ sec $0.00055$ sec 33 $1800$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $10$ $100$ $0.0004$ sec $0.0004$ sec $0.00057$ sec $10$ $500$ $0.00039$ sec $0.0004$ sec $0.00051$ sec $10$ $1000$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $10$ $1500$ $0.00041$ sec $0.00042$ sec $0.00055$ sec $10$ $1800$ $0.00041$ sec $0.00042$ sec $0.00052$ sec 2020 $100$ $0.0004$ sec $0.00042$ sec $0.0005$ sec 2020 $500$ $0.0004$ sec $0.00042$ sec $0.0005$ sec 2020 $1000$ $0.00042$ sec $0.00043$ sec $0.00052$ sec 2020 $1500$ $0.00043$ sec $0.00044$ sec $0.00054$ sec 2020 $1800$ $0.00044$ sec $0.00044$ sec $0.00054$ sec
 Figure 10: HPC Dataset vector embedding representation with addition of Noise
@@
 Secondly, we currently consider single-input and single-output operator modelling.
-Finally, our proposed NumData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples.
+Finally, our proposed NumTabData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples.
 This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available for a budget GPU).
@@
 In this paper, we presented a novel framework for the modelling of an analytic operator (such as a ML algorithm) when a large number of input data is available and thus no brute-force execution can be performed.
-We propose a deep learning model, NumData2Vec, which transforms a dataset to a lower $k$-dimensional representation space $z$.
-Our framework produces vector embeddings for the input datasets using NumData2Vec and performs a similarity search to identify the most relevant subset of datasets for any unseen input.
+We propose a deep learning model, NumTabData2Vec, which transforms a dataset to a lower $k$-dimensional representation space $z$.
+Our framework produces vector embeddings for the input datasets using NumTabData2Vec and performs a similarity search to identify the most relevant subset of datasets for any unseen input.
 By modelling the analytic operator based on this selected subset, we are able to accurately predict its output on any given input dataset.
 In practice, we demonstrated that our framework can accurately model various common algorithms and compared favourably against a similar recent framework [9], in both accuracy and speedup.
-Furthermore, we showed that NumData2Vec can create different vector representations for datasets from different scenarios.
-We also demonstrated that NumData2Vec can effectively detect when noise is introduced into a dataset.
+Furthermore, we showed that NumTabData2Vec can create different vector representations for datasets from different scenarios.
+We also demonstrated that NumTabData2Vec can effectively detect when noise is introduced into a dataset.
 {credits}